### v1

# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

Abstract
Achieving both realism and controllability in interactive closed-loop traffic simulation remains a key challenge in autonomous driving. Data-driven simulation methods reproduce realistic trajectories but suffer from covariate shift in closed-loop deployment, compounded by simplified dynamics models that further reduce reliability. Conversely, physics-based simulation methods enhance reliable and controllable closed-loop interactions but often lack expert demonstrations, compromising realism. To address these challenges, we introduce a dual-stage AV-centered simulation framework that conducts open-loop imitation learning pre-training in a data-driven simulator to capture trajectory-level realism and multimodality, followed by closed-loop reinforcement learning fine-tuning in a physics-based simulator to enhance controllability and mitigate covariate shift. In the fine-tuning stage, we propose RIFT, a simple yet effective closed-loop RL fine-tuning strategy that preserves the trajectory-level multimodality through a GRPO-style group-relative advantage formulation, while enhancing controllability and training stability by replacing KL regularization with the dual-clip mechanism. Extensive experiments demonstrate that RIFT significantly improves the realism and controllability of generated traffic scenarios, providing a robust platform for evaluating autonomous vehicle performance in diverse and interactive scenarios. Project Page: https://currychen77.github.io/RIFT/
1 Introduction
Reliable closed-loop traffic simulation is critical for developing advanced autonomous driving systems, supporting training and evaluation [1, 2]. An ideal simulation should possess two key properties: realistic, reflecting real-world driving behavior; controllable, enabling customizable interaction styles. Existing studies typically focus either on generating realistic traffic scenarios [3, 4] or on constructing controllable interactions [5, 6, 7]. Jointly ensuring realism and controllability in interactive closed-loop scenarios remains an open problem.
Traffic simulation approaches can be broadly categorized based on the underlying simulation platform, each presenting distinct trade-offs between realism and controllability, as illustrated in Figure 1. Methods based on data-driven simulators exploit real-world data to generate realistic trajectories by learning multimodal behavioral patterns through imitation learning (IL), conditioned on the initial scene context [8, 9]. However, their open-loop training paradigm introduces the covariate shift problem during closed-loop deployment, arising from the distribution mismatch between training and deployment states. Moreover, data-driven simulators often adopt a unified vehicle dynamics model for all agents and simplified environment dynamics [10, 11], resulting in unrealistic interactions and state transitions that further degrade closed-loop reliability. In contrast, physics-based simulators provide fine-grained control over scenario configuration through physical engines, enabling high-fidelity closed-loop interactions. Nonetheless, the absence of expert demonstrations makes it challenging to reproduce realistic behavior. To mitigate this, several approaches employ reinforcement learning (RL) to acquire controllable behaviors directly through interaction with the simulator [12, 13, 14, 15], though often at the cost of realism. Other approaches enhance realism by injecting real-world traffic data into physics-based simulators [16, 17], but typically rely on log-replay or rule-based simulation, limiting controllability and interactivity. Despite recent advances, a fundamental trade-off persists between realism and controllability across both paradigms, making it challenging to achieve both simultaneously in interactive closed-loop scenarios.
Figure 1: Traffic Simulation across Different Platforms. (a) Data-driven Simulator: employs imitation learning to replicate real-world driving behaviors, but suffers from covariate shift and simplified dynamics; (b) Physics-based Simulator: enables controllable scenario construction via high-fidelity closed-loop interaction, but lacks large-scale real-world data; (c) Our framework: combines open-loop IL pre-training in a data-driven simulator to ensure realism with closed-loop RL fine-tuning in a physics-based simulator to enhance controllability.
Drawing inspiration from the widely adopted “pre-training and fine-tuning” paradigm in large language models (LLMs) [18, 19, 20], we combine the strengths of two traffic simulation paradigms. Specifically, we perform open-loop IL pre-training in a data-driven simulator to capture trajectory-level realism and multimodality, followed by closed-loop RL fine-tuning in a physics-based simulator to address covariate shift and enhance controllability in interactive scenarios.
Building on this foundation, we introduce a dual-stage AV-centered simulation framework, as illustrated in Figure 1, which integrates the strengths of both data-driven and physics-based simulators through the “pre-training and fine-tuning” paradigm, thereby balancing realism and controllability in traffic simulation. Within this framework, critical background vehicles (CBVs) are first identified via route-level interaction analysis to construct interactive scenarios, focusing on agents most likely to interact with the AV. We then leverage an IL pre-trained trajectory generation model to produce multimodal and realistic CBV trajectories, ensuring trajectory-level realism. In the fine-tuning stage, we propose RIFT, a simple yet effective closed-loop RL fine-tuning strategy designed to enhance controllability and mitigate covariate shift. While previous approaches [21, 22] typically fine-tune only the best-performing trajectory or action, RIFT evaluates all candidate trajectories using a GRPO-style [20] group-relative advantage, preserving trajectory-level multimodality during fine-tuning. Additionally, by replacing KL regularization with a dual-clip surrogate objective, RIFT improves controllability and training stability in highly interactive closed-loop scenarios. Our contributions can be summarized as:
  • We propose a dual-stage AV-centered simulation framework that combines open-loop IL pre-training in a data-driven simulator to capture trajectory-level realism and multimodality with closed-loop RL fine-tuning in a physics-based simulator to enhance controllability and mitigate covariate shift. This decoupled paradigm leverages the complementary strengths of both platforms, balancing realism and controllability in interactive closed-loop scenarios.
  • We propose RIFT, a simple yet effective closed-loop RL fine-tuning strategy that enhances controllability and mitigates covariate shift in interactive closed-loop scenarios. During fine-tuning, RIFT maintains trajectory-level multimodality via a GRPO-style group-relative advantage formulation while enhancing controllability and training stability by replacing KL regularization with a dual-clip surrogate objective.
  • Extensive experiments demonstrate that RIFT enables the generation of realistic and controllable scenarios, and improves the reliability of AV evaluation under diverse and interactive scenarios.
2 Related Work
Realistic Traffic Simulation. A variety of generative architectures have been explored for realistic traffic simulation, including conditional variational autoencoders [23, 24, 25] and diffusion-based models [26, 27, 28]. However, maintaining long-term stability remains challenging due to the covariate shift between open-loop training and closed-loop deployment. Recent methods such as SMART [29], GUMP [30], Trajeglish [31], and MotionLM [32] address this issue by formulating traffic simulation as a next-token prediction (NTP) task, leveraging discrete action spaces to improve closed-loop robustness. Despite these advances, most approaches remain confined to data-driven simulation platforms [10, 11, 33, 34], which typically adopt a uniform vehicle dynamics model across agents and simplified environmental dynamics. Such oversimplifications limit the fidelity and reliability of long-term closed-loop interactions, especially in complex and interactive scenarios.
Controllable Traffic Simulation. To generate traffic scenarios aligned with user intent, recent studies have introduced diverse conditioning mechanisms. CTG [35] and MotionDiffuser [36] employ diffusion models conditioned on cost-based control signals. Language-conditioned methods, including CTG++[37], LCTGen[38], and ProSim [39], enable user specification through natural language prompts. Other strategies adopt guided sampling (SceneControl [40]), retrieval-based generation (RealGen [41]), or reward-driven causality modeling (CCDiff [42]). Though effective in improving controllability, these approaches are typically limited to open-loop or simplified environments, leaving the challenge of unifying realism and controllability in closed-loop simulation unresolved.
Closed-Loop Fine-Tuning. Covariate shift—the mismatch between open-loop training and closed-loop deployment—remains a key challenge for reliable long-term traffic simulation. To address this, recent work explores closed-loop fine-tuning strategies. Hybrid IL and RL methods [21, 22, 43] enhance robustness but typically fine-tune the entire model via RL, which often compromises realism due to the difficulty of designing human-aligned reward functions. Supervised fine-tuning approaches such as CAT-K [44] show strong performance but rely on expert demonstrations, limiting scalability. TrafficRLHF [45] improves alignment through reinforcement learning with human feedback (RLHF), but demands costly human input and suffers from reward model instability. Moreover, most existing methods focus on optimizing the best action or trajectory, ignoring the inherent multimodality of traffic simulation, thus limiting behavioral diversity during fine-tuning.
3 Background
3.1 Task Redefinition
Following the widely adopted paradigm for closed-loop evaluation and training in autonomous driving [46, 47], our simulation framework includes a single autonomous vehicle (AV) navigating a predefined global route, accompanied by multiple rule-based background vehicles (BVs), forming an AV-centered closed-loop simulation environment. These BVs either provide diverse interactive data for training or serve to evaluate the AV’s robustness. Building upon this setup, we identify a subset of critical background vehicles (CBVs) that exhibit a higher likelihood of interacting with the AV. For these CBVs, the rule-based control is replaced with a well-trained trajectory generation model, enabling the synthesis of realistic and controllable behaviors in interactive closed-loop scenarios.
3.2 CBV-Centered Realistic Trajectory Generation
With recent advances in imitation learning, data-driven approaches have demonstrated strong performance in generating realistic, multimodal trajectories [48, 49, 50, 51, 52]. In fully observable simulation environments, methods like Pluto [53] can generate more reliable trajectories by leveraging ground-truth states. Moreover, its modular encoder-decoder architecture also facilitates downstream fine-tuning. Based on these advantages, we adopt Pluto as our trajectory generation model for synthesizing realistic, multimodal trajectories. Implementation details are presented below.
CBV-Centered Scene Encoding. Following [53], for each CBV in the scene, we extract its current feature $F_{\mathrm{cbv}}$, the historical features of neighboring vehicles $F_{\mathrm{neighbor}}$, and vectorized map features $F_{\mathrm{map}}$. These features are encoded into $E_{\mathrm{cbv}}\in\mathbb{R}^{1\times D}$, $E_{\mathrm{neighbor}}\in\mathbb{R}^{N_{\mathrm{neighbor}}\times D}$, and $E_{\mathrm{map}}\in\mathbb{R}^{N_{\mathrm{map}}\times D}$, respectively, where $N_{\mathrm{neighbor}}$ and $N_{\mathrm{map}}$ denote the number of neighboring vehicles and map elements, and $D$ is the embedding dimension. To model the interactions among these embeddings, we concatenate them and apply a global positional embedding ($\mathrm{PE}$) to obtain the unified scene embedding $E_{s}\in\mathbb{R}^{(1+N_{\mathrm{neighbor}}+N_{\mathrm{map}})\times D}$ as:
$E_{s}$ $=\mathrm{concat}(E_{\mathrm{cbv}},E_{\mathrm{neighbor}},E_{\mathrm{map}})+\mathrm{PE}.$ (1)
This scene embedding $E_{s}$ is then passed through $N$ Transformer encoder blocks for feature aggregation, yielding the final CBV-centered scene embedding $E_{\mathrm{enc}}$. Each encoder block follows the standard Transformer formulation. Specifically, the $i$-th block is defined as:
$E^{i}_{s}=E^{i-1}_{s}+\mathrm{MHA}(\mathrm{LayerNorm}({E^{i-1}_{s}})),$ (2)
$E^{i}_{s}=E^{i}_{s}+\mathrm{FFN}\left(\mathrm{LayerNorm}(E^{i}_{s})\right),$
where $\mathrm{MHA}$ is the standard multi-head attention function, $\mathrm{FFN}$ is the feedforward network layer.
Multimodal Trajectory Decoding. To capture the multimodal nature of real-world driving behaviors, we adopt the longitudinal-lateral decoupling mechanism proposed in [53]. This approach leverages reference line information to construct high-level lateral queries $Q_{\mathrm{lat}}\in\mathbb{R}^{N_{\mathrm{ref}}\times D}$, and introduces learnable longitudinal queries $Q_{\mathrm{lon}}\in\mathbb{R}^{N_{\mathrm{lon}}\times D}$. These are concatenated and projected to form the multimodal navigation query $Q_{\mathrm{nav}}\in\mathbb{R}^{N_{\mathrm{ref}}\times{N_{\mathrm{lon}}}\times D}$ as:
$Q_{\mathrm{nav}}=\mathrm{Projection}(\mathrm{concat}(Q_{\mathrm{lat}},Q_{\mathrm{lon}})),$ (3)
where $N_{\mathrm{ref}}$ and $N_{\mathrm{lon}}$ denote the number of reference lines and longitudinal anchors, respectively. The navigation query $Q_{\mathrm{nav}}$ and the scene embedding $E_{\mathrm{enc}}$ are then fed into $N$ decoder blocks to model lateral, longitudinal, and cross-modal interactions. Each decoder block is structured as:
$\hat{Q}^{i-1}_{\mathrm{nav}}$ $=\mathrm{SelfAttn}(\mathrm{SelfAttn}(Q^{i-1}_{\mathrm{nav}},\mathrm{dim}=0),\mathrm{dim}=1),$ (4)
$Q^{i}_{\mathrm{nav}}$ $=\mathrm{CrossAttn}(\hat{Q}^{i-1}_{\mathrm{nav}},E_{\mathrm{enc}},E_{\mathrm{enc}}).$
$\mathrm{SelfAttn}$, $\mathrm{CrossAttn}$ denote multi-head self-attention and cross-attention operations, respectively. Given the decoder’s final output $Q_{\mathrm{dec}}$, two MLP heads are applied to produce the CBV-centered multimodal trajectories $\mathcal{T}\in\mathbb{R}^{N_{\mathrm{ref}}\times N_{\mathrm{lon}}\times T\times 6}$ and their corresponding confidence scores $\mathcal{S}\in\mathbb{R}^{N_{\mathrm{ref}}\times N_{\mathrm{lon}}}$:
${\mathcal{T}}=\mathrm{MLP}(Q_{\mathrm{dec}}),\;{\mathcal{S}}=\mathrm{MLP}(Q_{\mathrm{dec}}),$ (5)
where $T$ is the prediction horizon, and each trajectory point $\tau_{t}^{i}$ encodes $[p_{x},p_{y},\cos{\theta},\sin{\theta},v_{x},v_{y}]$.
4 Methodology
Leveraging the trajectory generation model described in Section 3.2, realistic and multimodal trajectories can be generated across diverse scenarios. However, the open-loop training paradigm leaves the policy vulnerable to covariate shift, a distribution mismatch between expert states observed during training and those encountered at deployment, even equipped with techniques such as contrastive learning [54, 55] or data augmentation [53]. To mitigate this issue, we propose RIFT, a simple yet effective closed-loop RL fine-tuning strategy designed to enhance controllability and mitigate covariate shift while maintaining trajectory-level realism and multimodality. The following sections detail RIFT’s implementation within the dual-stage AV-centered simulation framework.
Figure 2: Overview of the RIFT: The upper section illustrates the overall architecture of RIFT. To enhance controllability, only the trajectory scoring head is fine-tuned, with the rest of the pre-trained network kept frozen to preserve trajectory-level realism. (a) The CBV identification mechanism introduces route-level interactions between the AV and CBVs. (b) Closed-loop fine-tuning improves user-aligned controllability and mitigates covariate shift.
4.1 Critical Background Vehicle Identification
Following [1], we address the “curse of rarity” [56] by selectively intervening in a subset of critical background vehicles (CBVs) at key interaction moments, while retaining rule-based control for non-critical agents to reduce computational overhead.
To facilitate meaningful interactions between the AV and surrounding BVs, we adopt the principle of decoupling route-level interaction from behavior-level realism, as inspired by [57]. As illustrated in Figure 2, we estimate the route-level interaction probability between the AV’s predefined global route and the candidate routes of other BVs. The vehicle with the highest interaction probability is selected as the CBV, and its route is then synthesized based on the predicted conflict point to support further AV-centered closed-loop simulation. Further identification details are provided in Section B.2.
4.2 Closed-Loop RL Fine-Tuning
For the identified CBVs, we replace rule-based control with the IL pre-trained trajectory generation model (Section 3.2), which generates $N_{\mathrm{ref}}\times N_{\mathrm{lon}}$ candidate trajectories per CBV. The highest-scoring trajectory is selected for closed-loop execution. However, due to the covariate shift between open-loop training and closed-loop deployment, this approach often accumulates errors over time, resulting in unrealistic long-term behavior. To address this issue, prior work has explored RL [58, 59] and hybrid IL-RL approaches [21, 22] for closed-loop fine-tuning. However, these methods typically compute the loss only on the executed action or the best trajectory, overlooking the remaining candidates and neglecting the inherently multimodal nature of real-world driving behaviors. Motivated by the success of GRPO [20] in large language models fine-tuning, we propose to treat all the candidate trajectories $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N{\mathrm{ref}}\times N_{\mathrm{lon}}}$ as a group output. We then compute the corresponding trajectory-level group-relative reward =$\mathcal{R}=\{r_{i}\}_{i=1}^{N{\mathrm{ref}}\times N_{\mathrm{lon}}}$ and the group-relative advantage $\mathcal{A}=\{\hat{A}_{i}\}_{i=1}^{G}$ as follows:
$r_{i}=\sum_{t=0}^{T}\gamma^{t}\left[\mathrm{StateWiseRM}(\tau_{i}^{t})\right],\ \hat{A}_{i}=\frac{r_{i}-\mathrm{mean}(\mathcal{R})}{\mathrm{std}(\mathcal{R})},\quad\forall i=1,\dots,G$ (6)
where $G=N_{\mathrm{ref}}\times N_{\mathrm{lon}}$, and $\mathrm{StateWiseRM}$ is the user-defined state-wise reward model. With the group relative advantage $\mathcal{A}$, we derive the GRPO objective in our framework as:
$\begin{split}\mathcal{J}_{\mathrm{GRPO}}(\theta)&=\mathbb{E}_{[\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}(s)}]}\\ &\frac{1}{G}\sum_{i=1}^{G}\left\{\min\left[\frac{\pi_{\theta}(\tau_{i}|s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}\hat{A}_{i},\mathrm{clip}\left(\frac{\pi_{\theta}(\tau_{i}|s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)},1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right]-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\right\},\end{split}$ (7)
where ${\pi}_{\theta}$ is the current trajectory generation model, ${\pi}_{\mathrm{ref}}$ is the reference policy model, $\epsilon$ and $\beta$ are hyper-parameters, and $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}|s)/{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}$ is the trajectory-level importance sampling ratio.
By leveraging IL pre-training to capture trajectory-level realism and multimodality, our closed-loop RL fine-tuning serves as an alignment mechanism that bridges the gap between user preferences and real-world data distributions, while preserving trajectory-level realism and multimodality. Building on this foundation, we remove the KL regularization term used in the original GRPO formulation, as the IL pre-trained distribution may significantly diverge from the user-preferred output. In such cases, enforcing constraints via KL divergence is unnecessary and may hinder alignment.
Removing the KL regularization introduces instability during fine-tuning. PPO-Clip [58] addresses this by constraining updates through a clipped importance sampling ratio, effectively limiting policy shifts to a trust region. However, this mechanism is inadequate in our group-relative setting. The original clipping mechanism is designed for on-policy learning, where only the executed action is corrected. In contrast, we evaluate an entire set of trajectory candidates, where the trajectory-level importance ratio $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}|s)/{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}$ can become arbitrarily large, especially when $\pi_{\theta}(\tau_{i}|s)$ dominates $\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)$ and $\hat{A}_{i}<0$, the resulting product $\rho_{i}(\theta)\hat{A}_{i}$ yields large negative values, leading to high variance and unstable updates. To stabilize fine-tuning while preserving alignment with user-defined preferences, we adopt the dual-clip mechanism from Dual-Clip PPO [60, 61], where a constant $c>1$ acts as a lower bound on clipped negative advantages. The resulting RIFT objective is defined below (see Appendix A for theoretical analysis).
$\begin{split}\mathcal{J}_{\mathrm{RIFT}}(\theta)&=\mathbb{E}_{[\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}(s)}]}\\ &\frac{1}{G}\sum_{i=1}^{G}\left\{\max\left[\min\left[\frac{\pi_{\theta}(\tau_{i}|s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}\hat{A}_{i},\mathrm{clip}\left(\frac{\pi_{\theta}(\tau_{i}|s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)},1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right],c\hat{A}_{i}\right]\right\}.\end{split}$ (8)
5 Experiment
This section systematically investigates the following research questions: Q1: How do the realism and controllability of interactive scenarios generated by RIFT compare to the baselines? Q2: How can the generated scenarios be leveraged for downstream tasks in autonomous driving? Q3: What are the contributions of each component within RIFT to the overall performance, and how well is controllability preserved under different user-specified reward settings? We first present the experimental setup, including the environment, baseline methods, and evaluation metrics. We then analyze the results to address these research questions.
5.1 Experiment Setups
Under the dual-stage AV-centered simulation framework, we adopt Pluto [53] as our trajectory generation model, owing to its strong performance and publicly available implementation. To ensure fairness during pre-training, we directly utilize the IL pre-trained checkpoint released by Pluto, which was pre-trained on the nuPlan dataset [11]. This setup enables a focused investigation of closed-loop fine-tuning. We conduct simulations in CARLA [62], building upon the Bench2Drive [46] framework to support comprehensive AV-centered closed-loop simulation and evaluation. Additional details regarding implementation, framework design, training protocols, and evaluation settings are provided in Appendix B.
Baseline. To systematically evaluate the effectiveness of RIFT in closed-loop fine-tuning, we compare it against the following baselines (details are provided in Section B.5).
  • Pure RL/IL: Methods trained purely with either RL or IL without fine-tuning, including Pluto [53], FREA [14], FPPO-RS [14], and PPO [14].
  • RLFT/SFT: Methods that apply RL or supervised signals to fine-tune the pre-trained Pluto models, such as PPO-Pluto [58], REINFORCE-Pluto [59], GRPO-Pluto [20], and SFT-Pluto.
  • Hybrid: Methods that integrate both RL and supervised signals during fine-tuning, including RTR-Pluto [21] and RS-Pluto [22].
All methods are fine-tuned on the same module—the trajectory scoring head of Pluto—to ensure fair comparisons and decouple trajectory-level realism from behavior-level reliability, as confirmed by the ablation studies in Section 5.4. Following the realism standards of the Sim Agent Challenge in WOSAC [34], we adopt a normal reward setting for all RL-based baselines, with the detailed configuration provided in Section B.6. Results under an aggressive reward configuration are also reported in Section 5.4.
Metrics. We extend the WOSAC evaluation framework [34], which evaluates behavior controllability and realism across three key categories—motion, agent interactions, and map adherence—to define our own set of evaluation metrics, as inspired by [63]. More details are provided in Section B.7.
  • Controllability: We measure infractions using Scenario Collision Per Kilometer (CPK) and Off-Road Rate (ORR)[41, 42], evaluate driving progress via Route Progress (RP), and assess driving safety with safety-critical metrics such as 2D-TTC[64] and ACT [65].
  • Realism: Quantifying realism in traffic scenarios remains a challenging problem. Inspired by prior work that evaluates realism from a distributional perspective [21, 34, 63]. We use the Shapiro–Wilk test (SW) [66, 67] to assess the normality of speed and acceleration, and Wasserstein Distance (WD) [68] to measure deviation from the target speed in the CARLA [62].
Table 1: Comparison in Controllability and Realism. Metrics are evaluated under the PDM-Lite [69] AV setting across three random seeds, with the best and the second-best results highlighted accordingly.
Method Type Controllability Realism
Infraction Progress Safety Speed Acceleration
ORR $\downarrow$ CPK $\downarrow$ RP $\uparrow$ 2D-TTC $\uparrow$ ACT $\uparrow$ SW $\uparrow$ WD $\downarrow$ SW $\uparrow$
Pluto [53] IL 0.24 $\pm$ 0.15 5.06 $\pm$ 2.69 564.14 $\pm$ 114.41 2.50 $\pm$ 1.48 2.44 $\pm$ 1.39 0.88 $\pm$ 0.01 5.81 $\pm$ 0.06 0.90 $\pm$ 0.01
PPO [14] RL 9.17 $\pm$ 2.39 13.95 $\pm$ 2.34 409.51 $\pm$ 30.38 2.59 $\pm$ 1.60 2.52 $\pm$ 1.57 0.95 $\pm$ 0.01 4.45 $\pm$ 0.15 0.89 $\pm$ 0.02
FREA [14] RL 9.01 $\pm$ 2.09 30.42 $\pm$ 5.28 292.81 $\pm$ 68.54 2.71 $\pm$ 1.40 2.67 $\pm$ 1.41 0.93 $\pm$ 0.01 5.10 $\pm$ 0.14 0.93 $\pm$ 0.01
FPPO-RS [14] RL 8.60 $\pm$ 0.25 21.39 $\pm$ 3.23 356.79 $\pm$ 26.19 2.55 $\pm$ 1.69 2.53 $\pm$ 1.68 0.87 $\pm$ 0.01 5.80 $\pm$ 0.11 0.80 $\pm$ 0.03
SFT-Pluto SFT 0.06 $\pm$ 0.07 6.33 $\pm$ 2.23 780.48 $\pm$ 41.05 2.20 $\pm$ 1.64 2.12 $\pm$ 1.51 0.88 $\pm$ 0.02 6.01 $\pm$ 0.19 0.87 $\pm$ 0.02
RS-Pluto [22] SFT+RLFT 1.05 $\pm$ 0.31 4.11 $\pm$ 3.90 819.40 $\pm$ 74.07 2.27 $\pm$ 1.45 2.23 $\pm$ 1.43 0.93 $\pm$ 0.00 5.40 $\pm$ 0.15 0.92 $\pm$ 0.01
RTR-Pluto [21] SFT+RLFT 0.08 $\pm$ 0.09 6.98 $\pm$ 2.59 481.60 $\pm$ 70.19 2.55 $\pm$ 1.60 2.47 $\pm$ 1.51 0.85 $\pm$ 0.00 6.24 $\pm$ 0.16 0.81 $\pm$ 0.03
PPO-Pluto RLFT 0.07 $\pm$ 0.13 6.89 $\pm$ 3.19 683.57 $\pm$ 38.12 2.66 $\pm$ 1.50 2.60 $\pm$ 1.43 0.95 $\pm$ 0.01 4.96 $\pm$ 0.31 0.90 $\pm$ 0.02
REINFORCE-Pluto RLFT 1.37 $\pm$ 1.13 6.98 $\pm$ 0.86 813.70 $\pm$ 24.76 2.39 $\pm$ 1.64 2.30 $\pm$ 1.55 0.92 $\pm$ 0.01 5.63 $\pm$ 0.19 0.90 $\pm$ 0.02
GRPO-Pluto RLFT 0.10 $\pm$ 0.08 7.24 $\pm$ 4.04 892.65 $\pm$ 65.27 2.65 $\pm$ 1.44 2.61 $\pm$ 1.48 0.94 $\pm$ 0.04 4.96 $\pm$ 0.89 0.96 $\pm$ 0.00
RIFT-Pluto (ours) RLFT 0.36 $\pm$ 0.20 6.83 $\pm$ 2.62 995.33 $\pm$ 84.62 2.74 $\pm$ 1.30 2.71 $\pm$ 1.32 0.97 $\pm$ 0.01 4.46 $\pm$ 0.43 0.93 $\pm$ 0.01
Figure 3: Speed and Acceleration Distribution. RL-based methods produce interactive but often unnatural behaviors, while supervised methods typically lead to conservative behavior. In contrast, RIFT achieves higher interactivity, reflected by increased average speed and acceleration, and more realistic behavior from the distributional perspective, avoiding unnecessary hesitation while keeping safe interactions.
5.2 Realistic and Controllable Traffic Scenario Generation (Q1)
Main Results. To address Q1, we evaluate the controllability and realism of the generated scenarios across different CBV methods, with results summarized in Table 1. RIFT consistently outperforms all baselines in both aspects across most settings. While supervised learning methods achieve slightly lower infraction rates, this improvement is primarily due to their inherently conservative behavior, derived from the expert PDM-Lite [69], which prioritizes safety by avoiding risky maneuvers.
This conservative tendency is further highlighted in Figure 3, where supervised policies exhibit significantly lower speed and acceleration profiles. In contrast, RIFT strikes a more favorable balance between safety and interactivity. It achieves superior safety performance, as reflected by higher 2D-TTC and ACT scores, while avoiding the overly cautious behaviors typical of supervised approaches. As shown in Figure 3, RIFT demonstrates higher average speed and acceleration, indicating more interactive behavior, while maintaining realistic motion profiles.
Figure 4: Qualitative examples illustrating the robustness of RIFT across diverse AV-centered traffic scenarios. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
Figure 5: Temporal comparisons illustrating RIFT’s superior performance over other baselines under AV-centered closed-loop simulation. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
Qualitative Results. To further demonstrate the effectiveness of RIFT, we present simulation results across a range of traffic scenarios under diverse conditions, including varying weather (e.g., rain and fog) and traffic scenarios (e.g., merging and crossing), as illustrated in Figure 4. These results demonstrate the robustness and generalizability of RIFT in challenging environments. In addition, we compare closed-loop simulations against representative baselines, as shown in Figure 5. Baseline methods often suffer from trajectory instability or poor-quality generation in closed-loop settings. In contrast, RIFT consistently generates smooth, high-quality trajectories with better temporal consistency. Additional qualitative results are provided in Section D.3.
5.3 Realistic Traffic Scenarios for Closed-Loop AV Evaluation (Q2)
In addressing Q2, we investigate the integration of RIFT into downstream tasks, specifically in the closed-loop AV evaluation. Building on the closed-loop evaluation framework [46, 70] in CARLA, we replace its hand-crafted scenarios with traffic flows generated by various CBV methods while retaining the same global route settings. To ensure fair and consistent evaluation, we adopt the standard AV evaluation metrics and introduce a new metric, Blocked Rate (BR), to measure situations where CBVs obstruct the AV. We evaluate two publicly available AV algorithms: the rule-based PDM-Lite [69] and the learning-based PlanT [71]. The results, presented in Table 3, demonstrate that under the RIFT-generated scenarios, AVs achieve near-optimal Driving Score (DS), Route Completion (RC), and Infraction Score (IS), along with the lowest BR. These findings highlight RIFT’s effectiveness in minimizing blocking while maintaining realism and interactivity, underscoring its superiority in closed-loop AV evaluation. Further details are available in Appendix C.
Table 2: Comparison of AV Evaluation across CBV Methods. Each metric is evaluated across three random seeds, with the best and the second-best results highlighted accordingly.
Method PDM-Lite [69] PlanT [71]
DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ BR $\downarrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ BR $\downarrow$
Pluto [53] 77.84 $\pm$ 2.20 81.84 $\pm$ 2.20 0.96 $\pm$ 0.00 23.33 $\pm$ 5.77 42.52 $\pm$ 4.72 52.53 $\pm$ 5.29 0.79 $\pm$ 0.01 46.67 $\pm$ 15.28
PPO[14] 76.26 $\pm$ 0.12 76.26 $\pm$ 0.12 1.00 $\pm$ 0.00 30.00 $\pm$ 0.00 36.39 $\pm$ 1.11 58.30 $\pm$ 3.41 0.66 $\pm$ 0.05 43.33 $\pm$ 11.55
FREA[14] 83.53 $\pm$ 0.13 83.53 $\pm$ 0.13 1.00 $\pm$ 0.00 20.00 $\pm$ 0.00 39.61 $\pm$ 1.34 53.82 $\pm$ 2.20 0.73 $\pm$ 0.01 56.67 $\pm$ 5.77
FPPO-RS[14] 83.52 $\pm$ 0.09 83.52 $\pm$ 0.09 1.00 $\pm$ 0.00 20.00 $\pm$ 0.00 38.85 $\pm$ 4.91 52.27 $\pm$ 5.38 0.66 $\pm$ 0.07 50.00 $\pm$ 10.00
SFT-Pluto 86.09 $\pm$ 2.04 90.16 $\pm$ 1.82 0.97 $\pm$ 0.03 13.33 $\pm$ 5.77 39.41 $\pm$ 4.97 57.90 $\pm$ 3.72 0.70 $\pm$ 0.06 50.00 $\pm$ 0.00
RS-Pluto [22] 89.32 $\pm$ 1.41 89.32 $\pm$ 1.41 1.00 $\pm$ 0.00 13.33 $\pm$ 5.77 42.05 $\pm$ 4.08 59.58 $\pm$ 3.27 0.66 $\pm$ 0.02 50.00 $\pm$ 0.00
RTR-Pluto [21] 87.64 $\pm$ 1.56 91.64 $\pm$ 1.56 0.96 $\pm$ 0.00 10.00 $\pm$ 0.00 40.08 $\pm$ 2.38 55.87 $\pm$ 1.71 0.69 $\pm$ 0.04 50.00 $\pm$ 0.00
PPO-Pluto 85.63 $\pm$ 2.02 89.63 $\pm$ 2.02 0.96 $\pm$ 0.00 16.67 $\pm$ 5.77 41.86 $\pm$ 2.78 59.82 $\pm$ 2.90 0.66 $\pm$ 0.04 43.33 $\pm$ 5.77
REINFORCE-Pluto 92.17 $\pm$ 3.45 92.17 $\pm$ 3.45 1.00 $\pm$ 0.00 10.00 $\pm$ 10.00 45.25 $\pm$ 1.75 63.74 $\pm$ 5.91 0.67 $\pm$ 0.05 43.33 $\pm$ 5.77
GRPO-Pluto 89.86 $\pm$ 2.10 91.19 $\pm$ 3.46 0.99 $\pm$ 0.02 6.67 $\pm$ 5.77 47.24 $\pm$ 5.67 61.88 $\pm$ 0.71 0.75 $\pm$ 0.07 $3$$\pm$ 5.77
RIFT-Pluto (ours) 94.78 $\pm$ 1.37 94.78 $\pm$ 1.37 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 44.28 $\pm$ 3.15 61.94 $\pm$ 5.08 0.64 $\pm$ 0.02 40.00 $\pm$ 10.00
Table 3: Ablation study on RIFT. Evaluation under PDM-Lite [69] AV setting with three random seeds.
Method Controllability Realism
Infraction Progress Safety Speed Acceleration
ORR $\downarrow$ CPK $\downarrow$ RP $\uparrow$ 2D-TTC $\uparrow$ ACT $\uparrow$ SW $\uparrow$ WD $\downarrow$ SW $\uparrow$
w/ Aggressive 0.91 (+0.55) 8.41 (+1.58) 1053.76 (+58.43) 2.93 (+0.19) 2.88 (+0.17) 0.97 (+0.00) 3.89 (-0.57) 0.94 (+0.01)
w/ All-head 0.43 (+0.07) 7.84 (+1.01) 827.12 (168.21) 2.83 (+0.09) 2.76 (+0.05) 0.96 (-0.01) 4.70 (+0.24) 0.94 (+0.01)
w/ PPO-Clip 0.04 (-0.32) 2.03 (-4.80) 655.39 (-339.94) 2.57 (-0.17) 2.54 (-0.17) 0.91 (-0.06) 5.92 (+1.46) 0.94 (+0.01)
RIFT-Pluto (ours) 0.36 6.83 995.33 2.74 2.71 0.97 4.46 0.93
5.4 Ablation Study (Q3)
In exploring the design choices within RIFT, we conduct a series of ablation studies focusing on three key aspects: the choice of policy optimization clip (Dual-Clip vs. PPO-Clip), the fine-tuning module (Trajectory Scoring Head vs. All Head), and the reward design (Normal vs. Aggressive). All experiments are conducted under consistent settings, and results are summarized in Table 3.
Dual-Clip vs. PPO-Clip. We first evaluate the effect of the dual-clip mechanism by replacing it with the standard PPO-Clip (w/ PPO-Clip). This variant leads to more conservative and less realistic behaviors, as evidenced by reduced Progress and degraded Speed Realism metrics.
Trajectory Scoring Head vs. All Head. Next, we examine the impact of the fine-tuning module on performance by comparing RIFT fine-tuned using only the trajectory scoring head with a variant fine-tuned on all heads (w/ All Head). The latter results in a slight drop across Realism, Progress, and Infraction metrics, likely due to the disruption of trajectory-level realism captured during pre-training.
Normal vs. Aggressive. Finally, we explore the effect of different reward preferences on controllability by adopting a more aggressive reward focused on driving efficiency (w/ Aggressive). This variant improves Progress and Speed Realism but increases Infractions, reflecting the intended trade-off and confirming that RIFT allows flexible reward design while preserving user controllability. Additional qualitative results regarding controllability are presented in Appendix D.1.
6 Conclusion
In this work, we propose a dual-stage AV-centered simulation framework that conducts open-loop IL pre-training in a data-driven simulator to capture trajectory-level realism and multimodality, followed by closed-loop RL fine-tuning in a physics-based simulator to address the covariate shift problem and enhance controllability. During fine-tuning, we introduce RIFT, a simple yet effective closed-loop RL fine-tuning strategy that preserves the trajectory-level multimodality through a GRPO-style group-relative advantage formulation and enhances both controllability and training stability by replacing KL regularization with the dual-clip mechanism. Extensive experiments demonstrate that RIFT generates high-quality traffic scenarios that are both realistic and controllable, offering a reliable simulation environment for evaluating AV performance in diverse and interactive scenarios, thus bridging the gap between scenario generation and closed-loop AV evaluation.
Limitations and Future Work. Currently, RIFT adopts a simple state-wise reward model, $\mathrm{StateWiseRM}$, as detailed in Section B.6, to evaluate each trajectory point without forward simulation. This may result in inaccurate trajectory-level reward estimations over time, potentially affecting closed-loop RL fine-tuning performance (further discussed in Section D.2). Additionally, the current implementation of our traffic simulation framework supports only AV methods with fully observable, vectorized scene representations. Extending this framework to support end-to-end closed-loop training and evaluation remains a promising direction for future research.
Appendix
[…]
1
Appendix A Theoretical Analysis
This section thoroughly analyzes RIFT under a binary reward setting ($r\in\{0,1\}$) without any KL regularization. The dual-clip surrogate guarantees stable gradients and a globally convergent update in success-probability space.
A.1 Problem Setup
For a given state $s$, let $\pi_{\theta}(\tau\mid s)$ denote the probability of trajectory $\tau$ under policy parameters $\theta$. Define the success probability
$p_{\theta}(s)\;:=\;\Pr_{\tau\sim\pi_{\theta}(\cdot\mid s)}\bigl{(}r(\tau,s)=1\bigr{)},$
and write $p:=p_{\theta_{\text{old}}}(s)\in(0,1)$ for the current policy. Since $r\in\{0,1\}$, the mean and variance of the Bernoulli random variable are given by:
$\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}[r(\tau,s)]=p,\qquad\mathsf{Var}_{\tau\sim\pi_{\theta_{\text{old}}}}[r(\tau,s)]=p(1-p).$
The (group-)relative advantage [20] in Equation 6 therefore reduces to
$\hat{A}(\tau,s)=\begin{cases}\sqrt{\tfrac{1-p}{p}},&r(\tau,s)=1,\\[6.0pt] -\sqrt{\tfrac{p}{1-p}},&r(\tau,s)=0.\end{cases}$ (9)
Dual-Clip Surrogate. RIFT discards the KL term entirely and maximizes
$f_{\epsilon}^{\mathrm{RIFT}}(\rho_{i},\hat{A}_{i})=\max\!\Bigl{(}\min\!\bigl{(}\rho_{i}\hat{A}_{i},\,\operatorname{clip}(\rho_{i},1-\epsilon,1+\epsilon)\hat{A}_{i}\bigr{)},c\,\hat{A}_{i}\Bigr{)},$ (10)
where $\rho_{i}=\pi_{\theta}(\tau_{i}\mid s)/\pi_{\theta_{\text{old}}}(\tau_{i}\mid s)$ is the trajectory-level importance sampling ratio, $\epsilon>0$ is the clipping range, and $c>1$ caps the most negative value.
Remark A.1.
If $\hat{A}_{i}<0$ and $\rho_{i}\gg 1$, the naïve PPO/GRPO objective becomes a large negative value, producing unstable gradients. Because
$f_{\epsilon}^{\mathrm{RIFT}}(\rho_{i},\hat{A}_{i})\;\geq\;c\,\hat{A}_{i},$ (11)
the surrogate is never more negative than $c\hat{A}_{i}$, so
$\bigl{|}\nabla_{\theta}\mathbb{E}[f_{\epsilon}^{\mathrm{RIFT}}]\bigr{|}\;\leq\;\bigl{|}\nabla_{\theta}\mathbb{E}[\rho_{i}\hat{A}_{i}]\bigr{|}.$ (12)
Hence, dual-clip reduces the worst-case gradient magnitude, ensuring stability.
A.2 Iterative Update in the Success-Probability Space
Let $\theta_{\text{new}}\in\arg\max_{\theta}\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}\!\bigl{[}f_{\epsilon}^{\mathrm{RIFT}}(\rho_{i},\hat{A}_{i})\bigr{]}$ and denote the corresponding success probability by $p_{\text{new}}:=p_{\theta_{\text{new}}}(s)$. The update may therefore be viewed as a mapping
$p_{n}=h_{\epsilon}(p_{n-1}),$ (13)
with $p_{n-1}=p_{\theta_{\text{old}}}(s)$ and $p_{n}=p_{\text{new}}$. We prove:
  1. (i)
    $h_{\epsilon}:[0,1]\to(0,1)$ is well-defined and continuous;
  2. (ii)
    with suitable $(c,\epsilon)$ or a sufficiently small step size, $h_{\epsilon}$ is a contraction, yielding a unique fixed point $p^{\star}\in(0,1)$ and global convergence $p_{n}\to p^{\star}$.
A.3 Existence of a Fixed Point
Lemma A.2 (Continuity and Interval Mapping).
For every $p\in[0,1]$ the updated probability $h_{\epsilon}(p)$ lies in $(0,1)$, and $h_{\epsilon}$ is continuous on $[0,1]$.
Proposition A.3 (Existence).
At least one fixed point $p^{\star}\in(0,1)$ satisfies $p^{\star}=h_{\epsilon}(p^{\star})$.
A.4 Uniqueness and Global Convergence
Bounded derivatives. For $\hat{A}_{i}\geq 0$, $|\partial f_{\epsilon}^{\mathrm{RIFT}}/\partial\rho_{i}|\leq(1+\epsilon)|\hat{A}_{i}|;$ for $\hat{A}_{i}<0$ it is bounded by $|\hat{A}_{i}|$. Consequently, the dependence of $p_{\text{new}}$ on $p$ admits a global Lipschitz constant $L(c,\epsilon)$.
Contraction. If $L(c,\epsilon)<1$ or a step size $\eta<1/L(c,\epsilon)$ is used, $h_{\epsilon}$ is a strict contraction on $(0,1)$.
Theorem A.4 (Convergence of RIFT).
Assume $c>1$ and $\epsilon>0$. If $|h_{\epsilon}(p)-h_{\epsilon}(p^{\prime})|\leq\lambda|p-p^{\prime}|$ for some $\lambda<1$ and all $p,p^{\prime}\in(0,1)$—for example, by choosing a step size $\eta<1/L(c,\epsilon)$—then
  1. (a)
    $h_{\epsilon}$ possesses a unique fixed point $p^{\star}\in(0,1)$;
  2. (b)
    any sequence $p_{n}=h_{\epsilon}(p_{n-1})$ converges geometrically: $|p_{n}-p^{\star}|\leq\lambda^{n}|p_{0}-p^{\star}|$;
  3. (c)
    The dual-clip term eliminates gradient explosions caused by negative advantages, guaranteeing stable learning.
A.5 Detailed Proofs
Throughout this subsection, we reuse the notation introduced in Section A.4:
$A^{+}(p)=\sqrt{\frac{1-p}{p}},\quad A^{-}(p)=-\sqrt{\frac{p}{1-p}},\quad\rho^{+}=1+\epsilon,\;\rho^{-}=1-\epsilon,\quad p\in(0,1),\ \epsilon>0,\ c>1.$
Proof of Lemma A.2.
The proof consists of three parts.
Boundary points. If $p=0$, every trajectory under the old policy fails and $A^{+}(0^{+})=\infty$. For any candidate new policy with success probability $q\in(0,1]$,
$\mathcal{J}(q)=q\,\rho^{+}A^{+}(0^{+})+(1-q)\,c\,A^{-}(0^{+})=(1+\epsilon)\,q\cdot\infty>0,\qquad\mathcal{J}(0)=0.$
Hence $h_{\epsilon}(0)=\arg\max_{q}\mathcal{J}(q)>0$. The case $p=1$ is symmetric, yielding $h_{\epsilon}(1)<1$.
Interior points. Let $p\!\in\!(0,1)$ and denote the step in parameter space by $\Delta\theta$. Smoothness of $\theta\!\mapsto\!\pi_{\theta}$ implies $\bigl{|}\log\rho_{i}\bigr{|}\!\leq\!K\|\Delta\theta\|$ for some $K<\infty$. Because $f_{\epsilon}^{\mathrm{RIFT}}$ is $(1+\epsilon)$-Lipschitz in $\rho_{i}$, the expected surrogate, cannot change by more than $K(1+\epsilon)\|\Delta\theta\|$ in one update; choosing $\|\Delta\theta\|<\!(K(1+\epsilon))^{-1}$ keeps $h_{\epsilon}(p)$ strictly inside $(0,1)$.
Continuity. Set $\Psi(\theta,p)=\mathbb{E}_{\tau\sim\pi_{\theta}}\![f_{\epsilon}^{\mathrm{RIFT}}]$. Joint continuity of $\Psi$ and compactness of the trajectory space (after clipping) satisfies the hypotheses of Berge’s maximum theorem, so the maximizer correspondence $\arg\max_{\theta}\Psi(\theta,p)$ is upper-hemicontinuous; the composition yields continuity of $h_{\epsilon}$ on $[0,1]$. ∎
Proof of Proposition A.3.
Define $g(p)=h_{\epsilon}(p)-p$. Lemma A.2 gives $g(0)>0$ and $g(1)<0$, while Lemma A.2 also guarantees the continuity of $g$. The Intermediate Value Theorem therefore yields a root $p^{\star}\in(0,1)$ with $p^{\star}=h_{\epsilon}(p^{\star})$. ∎
Proof of Theorem A.4.
The proof proceeds in four steps.
Step 1 – maximized surrogate. Substituting $\rho^{+},\rho^{-}$ into the surrogate gives
$\Phi(p)=p\,\rho^{+}A^{+}(p)+(1-p)\,\rho^{-}A^{-}(p)=2\epsilon\,\sqrt{p(1-p)}.$ (14)
Step 2 – Lipschitz bound.
$\Phi^{\prime}(p)=\epsilon\,\frac{1-2p}{\sqrt{p(1-p)}},\quad|\Phi^{\prime}(p)|\leq\frac{\epsilon}{\sqrt{p(1-p)}}.$ (15)
Fix $\delta\in(0,\tfrac{1}{2})$ such that the first iterate satisfies $h_{\epsilon}(p_{0})\in[\delta,1-\delta]$ (Lemma A.2 ensures such a $\delta$ exists). Then $|\Phi^{\prime}(p)|\leq L_{\delta}(\epsilon):=\epsilon/\sqrt{\delta(1-\delta)}$ for every subsequent $p$.
Step 3 – contraction. Using a gradient step $p_{n}=p_{n-1}+\eta\,\Phi^{\prime}(p_{n-1})$ with $\eta\in(0,1/L_{\delta}(\epsilon))$ gives
$|p_{n}-p_{n-1}|\leq\eta\,L_{\delta}(\epsilon)\,|p_{n-1}-p_{n-2}|=:\lambda\,|p_{n-1}-p_{n-2}|,\quad\lambda<1.$ (16)
Thus $h_{\epsilon}$ is a strict contraction on $([\delta,1-\delta],|\cdot|)$.
Step 4 – Banach fixed-point theorem. Because the metric space $\bigl{(}[\delta,1-\delta],\,|\cdot|\bigr{)}$ is complete and $h_{\epsilon}$ is a strict contraction on it, Banach’s fixed-point theorem guarantees
  1. (a)
    the existence and uniqueness of a fixed point $p^{\star}\in(\delta,1-\delta)\subset(0,1)$ such that $p^{\star}=h_{\epsilon}(p^{\star})$;
  2. (b)
    geometric convergence of the iterates generated by $p_{n}=h_{\epsilon}(p_{n-1})$:
    $|p_{n}-p^{\star}|\;\leq\;\lambda^{\,n}\,|p_{0}-p^{\star}|,\qquad 0<\lambda<1.$ (17)
Step 5 – bounded negative-advantage gradients. For every $\hat{A}_{i}<0$
$f_{\epsilon}^{\mathrm{RIFT}}(\rho_{i},\hat{A}_{i})=\max\!\bigl{(}\min(\rho_{i}\hat{A}_{i},\rho^{-}\hat{A}_{i}),\,c\hat{A}_{i}\bigr{)}\geq c\hat{A}_{i},$ (18)
so $|\partial f_{\epsilon}^{\mathrm{RIFT}}/\partial\rho_{i}|\leq|\hat{A}_{i}|$, preventing gradient explosions and guaranteeing stability. ∎
Appendix B Experimental Details
B.1 Experiment Framework
To enable reliable AV-centered closed-loop simulation, we develop our framework based on widely adopted closed-loop traffic simulation platforms, such as the CARLA Leaderboard [70] and Bench2Drive [46]. Traditionally, these platforms use predefined scenarios along the AV’s global route to evaluate the multi-dimensional performance of AV methods. In contrast, we replace these static scenarios with dynamically generated traffic flows by randomly spawning background vehicles around the AV’s global path and simulating their behavior using rule-based driving policies, as described in Section 3.1. Through the CBV identification mechanism outlined in Section B.2, we naturally introduce interactions between the AV and CBVs, thereby generating continuous, interactive scenarios over time. This framework serves as the foundation for both the training and evaluation processes in this paper.
B.2 CBV Identification Details
Identifying Critical Background Vehicles (CBVs) is essential to our AV-centered closed-loop simulation. Let $\mathcal{V}_{\text{AV}}$ denote the autonomous vehicle (AV), and $\mathcal{V}_{\text{BV}}=\{\mathcal{V}_{i}\}_{i=1}^{N}$ represent the set of background vehicles in the environment. The AV navigates along a predefined global route $\mathcal{P}=\{p_{k}\}_{k=1}^{M}$, where each $p_{k}$ corresponds to a waypoint along the route. The goal of CBV identification is to select background vehicles that are likely to share the AV’s destination and have similar estimated travel distance, thereby facilitating route-level interactions between the AV and CBVs. The primary criterion for identifying CBVs is the relative distance-to-goal difference between the AV and each background vehicle. This is mathematically expressed as:
$\left|\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{i})-\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{\text{AV}})\right|<\delta,$ (19)
where, $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{i})$ and $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{\text{AV}})$ denote the estimated travel distance required for the background vehicle $\mathcal{V}_{i}$ and the AV to reach waypoint $p_{k}$, respectively. The distance-to-goal for each vehicle is computed by determining the distance from its current position to the target waypoint $p_{k}$ using the A* global path planning algorithm [72]. A threshold $\delta$ is introduced to define the maximum allowable difference in distance-to-goal. A background vehicle is considered critical and included in the CBV set $\mathcal{C}$ if the absolute distance-to-goal difference between it and the AV is smaller than $\delta$.
This approach selects background vehicles whose destinations and estimated travel distances are sufficiently aligned with those of the AV, thereby ensuring meaningful and realistic route-level interactions. Once a CBV is identified, the planning path previously generated via A* during distance-to-goal estimation is directly adopted as its global navigation path, naturally introducing route-level interactions between the AV and CBVs. The threshold $\delta$ serves as a tunable parameter to adjust the sensitivity of the CBV selection process. In this study, we set $\delta$ to $15m$ to achieve a balanced trade-off between sensitivity and selection accuracy.
B.3 Algorithm Framework
For clarity, we summarize the procedure of RIFT within our AV-centered closed-loop simulation framework in Algorithm 1. The trajectory generation model initializes from the IL pre-trained checkpoint provided by Pluto official codebase [^1] https://github.com/jchengai/pluto, followed by closed-loop RL fine-tuning within the CARLA simulator [62] to generate realistic and controllable traffic scenarios.
Algorithm 1 Procedure for RIFT in the AV-Centered Closed-Loop Simulation Framework.
1:Input: IL pre-trained trajectory generation model $\pi_{\theta_{\text{init}}}$, buffer $\mathcal{D}$triangleright IL pre-training (nuPlan [11])
2:trajectory generation model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$
3:for $\text{iteraction}=1,\dots,I$ do $\triangleright$ Closed-loop RL fine-tuning (CARLA [62])
4:     Update the old trajectory generation model $\pi_{\theta_{old}}\leftarrow\pi_{\theta}$
5:     while $\mathcal{D}$ not full do $\triangleright$ Collect rollout data
6:         for $\text{step}=1,\dots,T$ do
7:
Generate $G$ trajectories $\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}$ for each CBV $\triangleright$ Policy inference
8:
Compute reward $\{r_{i}\}_{i=1}^{G}$, advantage $\{\hat{A}_{i}\}_{i=1}^{G}$ for each trajectory $\tau_{i}$ with Equation 6
9:
Store transition into buffer $\mathcal{D}$
10:         end for
11:     end while
12:     for RIFT $\text{iteration}=1,\dots,\mu$ do $\triangleright$ Policy fine-tuning
13:
Sample mini-batches transition from the buffer $\mathcal{D}$
14:
Update model $\pi_{\theta}$ by maximizing the RIFT objective (Equation 8)
15:     end for
16:end for
17:Output: RL fine-tuned trajectory generation model
B.4 Training Details
We perform closed-loop fine-tuning on selected modules of the IL pre-trained trajectory generation model (Pluto). As shown in the ablation results (Section 5.4), fine-tuning only the trajectory scoring head achieves the best trade-off between realism and controllability. Accordingly, all fine-tuning baselines adopt this setting to ensure consistency and fair comparison. Our training framework is built on the open-source Lightning platform [^2] https://github.com/Lightning-AI/pytorch-lightning. Fine-tuning is conducted on 2×2$\times$2 × Bench2Drive220, while evaluation is performed on dev10, both from the Bench2Drive project. All experiments are conducted on NVIDIA GeForce RTX 4090D GPUs, with each fine-tuning run taking approximately 8 hours on a single GPU. Detailed training setups and hyperparameter configurations are provided in Table 6 and Table 6.
B.5 Baselines Detailed Description
To comprehensively evaluate RIFT in an AV-centered closed-loop simulation environment, we compare it against a range of baselines, including pure imitation learning (IL), pure reinforcement learning (RL), and various fine-tuning approaches based on IL, RL, or their combination. We initialize all fine-tuning methods from the pre-trained Pluto checkpoint and fine-tune only the trajectory scoring head to preserve trajectory-level realism. The details of each baseline are summarized below.
  • Pluto [53] is an open-source IL-based planning framework for autonomous driving. It processes vectorized scene representations as input and outputs multimodal trajectories for downstream planning. In AV-centered closed-loop simulation, the method directly uses a pre-trained checkpoint without additional fine-tuning.
  • FREA [14] is an RL-based approach designed to generate safety-critical yet AV-feasible scenarios. It incorporates a feasibility-aware training objective. In the AV-centered closed-loop simulation, FREA selects potential collision points along the AV’s global route as adversarial goals.
  • PPO [14] is a variant of FREA that focuses solely on generating safety-critical scenarios. Unlike FREA, it disregards the feasibility constraints of AV and treats adversariality as the only optimization objective.
  • FPPO-RS [14] is another FREA variant that integrates AV’s feasibility constraints into the reward shaping process, thereby balancing adversariality with scenario reasonability.
  • PPO-Pluto fine-tunes the pre-trained trajectory generation model using the PPO algorithm [58]. The fine-tuning follows the same reward structure as detailed in Section B.6, aligning with RIFT.
  • REINFORCE-Pluto employs the REINFORCE algorithm [59] to fine-tune the pre-trained Pluto model under the same reward design as detailed in Section B.6.
  • GRPO-Pluto utilizes the basic GRPO algorithm [20] for fine-tuning, employing the pre-trained Pluto model as the reference for KL regularization, while incorporating the standard PPO-Clip.
  • SFT-Pluto is a purely supervised fine-tuning approach, where PDM-Lite [69] serves as the expert model, providing supervision at the target speed level.
  • RTR-Pluto [21] is a hybrid framework combining imitation and reinforcement learning. While the original RTR utilizes human driving trajectories as supervision, our setting replaces this with PDM-Lite due to the lack of human-level demonstrations. The RL component uses sparse infraction-based rewards, consistent with the original RTR, and applies PPO for optimization.
  • RS-Pluto [22] also adopts a hybrid IL+RL paradigm, originally trained via REINFORCE using ground-truth supervision and sparse rewards to ensure safety and realism. In our adaptation, PDM-Lite substitutes the ground-truth expert, while the rest of the methodology remains unchanged.
B.6 State-Wise Reward Model Setup
To capture diverse human driving styles, we decompose driving behaviors into distinct reward components, following [73]. Different styles are constructed by combining weights assigned to each reward component (detailed in Table 6), enabling a range of behaviors from aggressive to conservative. The total driving reward is defined as:
$R=R_{\text{collision}}+R_{\text{off-road}}+R_{\text{comfort}}+R_{\text{lane}}+R_{\text{velocity}}+R_{\text{timestep}}.$ (20)
The individual terms are described as follows:
  • $R_{\text{collision}}=-\left(\alpha_{\text{collision}}+\lvert v\rvert\right)\mathbbm{1}_{\mathrm{collision}}$: penalizes collisions, with higher penalties at higher speeds.
  • $R_{\text{off-road}}=-\alpha_{\text{boundary}}\mathbbm{1}_{\text{boundary}}$ $\alpha_{\text{boundary}}$: penalizes deviations from the drivable area.
  • $R_{\text{comfort}}=-\alpha_{\text{comfort}}\left(\mathbbm{1}_{\lvert a\rvert>4}+\mathbbm{1}_{\lvert\omega\rvert>4}\right)$ penalizes excessive acceleration and angular acceleration.
  • $R_{\text{l-align}}=\alpha_{\text{l-align}}\left(\min\left(\cos\left(\theta_{f}\right),0\right)+\alpha_{\text{vel-align}}\min\left(\cos\left(\theta_{f}\right)*v,0\right)+0.25\left(1-\frac{\lvert\theta_{f}\rvert}{\pi/2}\right)\right)$: guides the agent to follow the correct driving direction and remain parallel to the lane markings.
  • $R_{\text{l-center}}=-\alpha_{\text{l-center}}\left(\mathbbm{1}_{\cos\left(\theta_{f}\right)>0.5}*\left(\lvert x_{f}-\alpha_{\text{center-bias}}\rvert-\frac{0.05}{\exp\left(\lvert x_{f}-\alpha_{\text{center-bias}}\rvert-0.5\right)}\right)\right)$: guides the agent to prefer trajectories that remain centered within the lane.
  • $R_{\text{velocity}}=\alpha_{\text{velocity}}\max\left(\cos\left(\theta_{f}\right),0.0\right)\,\mathbbm{1}_{3<\lvert v\rvert<20}*\lvert v\rvert$: promotes forward movement and biases the agent toward choosing routes with consistent traffic flow rather than traffic jams.
  • $R_{\text{timestep}}=-\alpha_{\text{timestep}}\,\mathbbm{1}_{\lvert v\rvert>0\,\lor\,\lvert a\rvert>0}$ applies a small per-step penalty, encouraging efficiency. It is disabled when the agent is stationary to allow appropriate waiting behavior at intersections.
Building on the reward definitions above, we construct a state-wise reward model $\mathrm{StateWiseRM}(\cdot)$, which computes a scalar reward based on a set of interpretable features extracted from each trajectory point $\tau^{t}_{i}$. Specifically, we define a feature extraction function $\phi(\tau^{t}_{i})$ as:
$\phi(\tau^{t}_{i})=(\mathbbm{1}_{\text{collision}},\mathbbm{1}_{\text{boundary}},a_{\text{long}},a_{\text{lat}},\theta_{f},x_{f},v,a),$ (21)
where:
  • $\mathbbm{1}_{\text{collision}}$ and $\mathbbm{1}_{\text{boundary}}$ are binary indicators of potential collisions and off-road violations;
  • $a_{\text{long}}$ and $a_{\text{lat}}$ denote the longitudinal and lateral acceleration;
  • $v$ and $a$ are the magnitudes of velocity and acceleration;
  • $x_{f}$ is the lateral distance to the nearest lane centerline;
  • $\theta_{f}$ is the heading deviation with respect to the lane direction.
The state-wise reward is then computed as:
$r^{t}_{i}=\mathrm{StateWiseRM}(\phi(\tau^{t}_{i})).$ (22)
All features, except the infraction indicators, are directly derived from the trajectory. To estimate future infractions, we follow the forecasting model in [69] to simulate other agents’ future positions based on current states and actions and identify collisions via bounding box overlap. Off-road violations are detected using the Euclidean Signed Distance Field (ESDF) [53].
Table 4: Hyperparameters used in RIFT Training.
Parameter Value
Batch size 256
Rollout buffer capacity 4096
Fine-tune initial LR $1\times e^{-4}$
Minimum LR $1\times e^{-6}$
LR decay across iteration 0.9
LR schedule Cosine
Num. RIFT epoch 16
Warmup Epoch of RIFT 3
AdamW weight-decay $1\times e^{-5}$
Table 5: Hyperparameters of RIFT or RL baselines.
Parameter Value
PPO clipping ratio $\epsilon$ 0.2
Dual-clip ratio $c$ 3
Discount factor $\gamma$ 0.98
$\lambda_{\text{GAE}}$ [74] 0.98
Hidden dimension $D$ 128
Num. lon. queries $N_{lon}$ 12
Traj. time horizon $T$ 80
Map radius 120m
Frame rate 10Hz
Table 6: Reward Parameters for Different Driving Styles.
Parameter Normal Aggressive
$\alpha_{\text{collision}}$ 20.0 5.0
$\alpha_{\text{boundary}}$ 5.0 5.0
$\alpha_{\text{comfort}}$ 0.8 0.8
$\alpha_{\text{l-align}}$ 0.5 0.5
$\alpha_{\text{vel-align}}$ 0.05 0.05
$\alpha_{\text{l-center}}$ 0.6 0.6
$\alpha_{\text{center-bias}}$ 0.0 0.0
$\alpha_{\text{velocity}}$ 0.1 0.2
$\alpha_{\text{timestep}}$ 0.1 0.1
B.7 Controllability and Realism Metrics
Controllability. Following the metric design principles proposed in the WOSAC challenge [34] and other widely adopted evaluation frameworks [42, 63], we adopt a set of well-established metrics to comprehensively assess different aspects of controllability:
  • Scenario Collision Per Kilometer (CPK) [63]: the average number of scenario collisions per kilometer of driving distance.
  • Off-Road Rate (ORR) [63]: the percentage of time that CBVs spend off-road on average.
  • Route Progress (RP) [63]: the total distance traveled by all CBVs, reflecting route completion.
  • 2D Time-to-Collision (2D-TTC) [64]: the minimum of longitudinal and lateral time-to-collision from the AV’s perspective, capturing the immediate interaction risk posed by CBVs.
  • Anticipated Collision Time (ACT) [65] : a safety-critical metric measuring the AV’s proximity to potential collisions, reflecting the interaction intensity introduced by CBVs.
Realism. Following [63], most realism metrics rely on ground-truth trajectories of human drivers to evaluate how closely simulation agents mimic real-world behavior. However, since all our experiments are conducted in the CARLA simulator without expert demonstrations, we adopt distribution-level metrics to assess the realism of CBV behavior in terms of speed and acceleration. Specifically, we employ the following two metrics:
  • Wasserstein Distance (WD) [68]: measures the distance between two distributions $\mu$ and $\nu$. Since CARLA provides a predefined target speed for agents, we use WD to compare the simulated CBV speed distribution with the target speed distribution as the reference.
    $\text{WD}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\mathbb{E}_{(x,y)\sim\gamma}\left[d(x,y)\right].$ (23)
  • Shapiro–Wilk test (SW) [66]: assesses whether the distribution of speed and acceleration follows a normal distribution, capturing the statistical naturalness of CBV motion.
    $\text{SW}=\frac{\left(\sum_{i=1}^{n}a_{i}x_{(i)}\right)^{2}}{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}},$ (24)
    where $a_{i}$ are coefficients, $x_{(i)}$ are the ordered data points, $x_{i}$ are the sample values, $\bar{x}$ is the sample mean, and $n$ is the number of data points.
Appendix C AV Evaluation Details
C.1 AV Methods Implementation
To assess the effectiveness of RIFT in generating reliable and interactive scenarios for AV evaluation in the AV-centered closed-loop simulation environment, we evaluate the following representative and stable AV methods:
  • PDM-Lite [69]: A rule-based privileged expert method that achieves state-of-the-art performance on the CARLA Leaderboard 2.0 by leveraging components such as the Intelligent Driver Model and the kinematic bicycle model. This open-source method serves as a strong baseline for comparison.
  • PlanT [71]: An explainable, learning-based planning method that operates on an object-level input representation and is trained through imitation learning.
C.2 AV Evaluation Metrics
As detailed in Sections B.1 and B.4, we develop an AV-centered closed-loop simulation environment, including a training and evaluation pipeline based on Bench2Drive. The AV closed-loop evaluation metrics proposed in Bench2Drive extend the original metrics of the CARLA Leaderboard by emphasizing the specific strengths and weaknesses of different methods across various aspects, such as merging and overtaking, thereby making them suitable for evaluating performance under predefined scenarios. However, as noted in Section B.1, replacing predefined scenarios with CBV-generated traffic flows precludes the evaluation of specific AV capabilities. Since our primary objective is to validate the effectiveness of RIFT in providing reliable traffic scenarios for AV evaluation, we follow the default evaluation metrics from CARLA Leaderboard to assess the overall driving performance of the AV. Additionally, we introduce a new metric to quantify the AV blockage events, which may arise when traffic flows are unreliable. The evaluation metrics are summarized as follows:
  • Driving Score (DS): $R_{i}P_{i}$ — The main metric of the leaderboard, calculated as the product of route completion and the infraction penalty. Here, $R_{i}$ represents the percentage of completion of the $i$-th route, and $P_{i}$ denotes the infraction penalty. The maximum value is 100.
  • Route Completion (RC): $R_{i}$ — The percentage of the route distance completed by the agent. The maximum value is 100.
  • Infraction penalty (IP): $P_{i}=\prod_{j}p_{j}^{\text{infractions}_{j}}$ — This metric tracks various types of infractions and aggregates all instances triggered by the agent as a geometric series. The agent starts with an ideal base score of 1.0, which decreases with each infraction to a minimum of 0.
  • Block Rate (BR): The average number of occurrences where a CBV fails to navigate its route normally and obstructs the AV’s progress.
Appendix D Additional Results
D.1 Detailed Qualitative Results of Controllability
As discussed in Section 5.4, we investigate the controllability of RIFT under different reward configurations. The aggressive variant applies a reduced collision penalty and places greater emphasis on driving efficiency (Table 6), encouraging assertive behaviors such as overtaking. In contrast, the normal configuration imposes a higher collision penalty to promote safer and more conservative driving behaviors.
Quantitative results in Table 3 show that the aggressive variant achieves higher Progress and Speed Realism scores, albeit at the cost of increased infractions. To complement these findings, Figure 6 presents a qualitative comparison in a single-lane intersection scenario where a leading BV halts at a stop sign. The aggressive CBV variant attempts an overtaking maneuver, resulting in a collision, whereas the normal CBV variant yields and waits, demonstrating distinct behavioral patterns induced by different reward preferences. These results highlight the controllability of RIFT in modulating driving behaviors according to user-specified reward configurations.
Figure 6: Qualitative illustration of RIFT’s controllability under different reward configurations. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
D.2 Detailed Analysis in Driving Comfort
Metrics. To further evaluate the driving comfort of different CBV methods, we define several comfort metrics based on Bench2Drive, which assesses agent comfort through acceleration and jerk profiles. Specifically, we measure comfort using the following metrics:
  • Uncomfortable Rate (UCR): the percentage of simulation time during which CBVs experience discomfort.
  • Driving Jerk (Jerk): the time derivative of acceleration, quantifying the abruptness of acceleration changes and the smoothness of CBV trajectories.
To determine whether a CBV’s current state is considered comfortable, we adopt the Frame Variable Smoothness (FVS) criterion from Bench2Drive:
$\begin{split}\text{Frame Variable Smoothness (FVS)}&=\begin{cases}\text{True}&\text{if lower bound}\leq p_{\,i}\leq\text{upper bound}.\\ \text{False}&\text{otherwise}\end{cases}\\ &p\in\text{smoothness vars},\,0\leq i\leq\text{total frames}\end{split}$ (25)
The smoothness variables include longitudinal acceleration (expert bounds: [-4.05, 2.40]), maximum absolute lateral acceleration (expert bounds: [-4.89, 4.89]), and maximum jerk magnitude (expert bounds: [-8.37, 8.37]).
Table 7: Comparison of CBV Comfort Metrics across various AV Methods. Each metric is evaluated across three random seeds.
Method PDM-Lite [69] PlanT [71]
UCR $\downarrow$ Jerk $\downarrow$ UCR $\downarrow$ Jerk $\downarrow$
Pluto [53] 56.45 $\pm$ 4.14 -0.16 $\pm$ 3.72 50.26 $\pm$ 2.17 -0.42 $\pm$ 3.38
PPO [14] 74.76 $\pm$ 2.71 -0.51 $\pm$ 4.61 74.90 $\pm$ 1.21 0.40 $\pm$ 4.83
FREA [14] 72.40 $\pm$ 1.72 0.29 $\pm$ 4.61 73.48 $\pm$ 3.83 -0.15 $\pm$ 4.91
FPPO-RS [14] 68.33 $\pm$ 1.90 -0.07 $\pm$ 3.96 66.67 $\pm$ 0.82 -0.15 $\pm$ 3.95
SFT-Pluto 68.14 $\pm$ 4.91 -0.06 $\pm$ 4.06 59.78 $\pm$ 4.72 -0.11 $\pm$ 4.00
RS-Pluto [22] 70.31 $\pm$ 4.07 0.32 $\pm$ 4.12 65.18 $\pm$ 2.11 -0.16 $\pm$ 4.07
RTR-Pluto [21] 55.58 $\pm$ 4.76 -0.19 $\pm$ 3.37 45.12 $\pm$ 2.66 -0.14 $\pm$ 3.34
PPO-Pluto 58.29 $\pm$ 2.70 -0.32 $\pm$ 3.70 54.85 $\pm$ 5.82 -0.07 $\pm$ 3.40
REINFORCE-Pluto 68.10 $\pm$ 1.22 0.23 $\pm$ 3.96 64.94 $\pm$ 5.36 -0.11 $\pm$ 3.96
GRPO-Pluto 78.58 $\pm$ 0.59 0.22 $\pm$ 4.62 77.13 $\pm$ 0.65 -0.23 $\pm$ 4.58
RIFT-Pluto (ours) 76.90 $\pm$ 2.82 0.59 $\pm$ 4.12 72.41 $\pm$ 4.02 0.21 $\pm$ 4.44
Figure 7: Controller Performance.
Main Results. The quantitative results of the comfort metrics are presented in Table 7. All CBV methods exhibit notable levels of driving discomfort. Although the more conservative methods identified in Section 5.2 achieve relatively lower discomfort levels, a high baseline of discomfort persists across methods.
To investigate the underlying causes of discomfort, we further decouple the planned trajectories from the executed control actions. In CARLA, most CBV methods rely on PID controllers to transform high-level trajectory waypoints into executable driving commands, including throttle, steering, and brake. As shown in Figure 7, the upper panel illustrates the speed tracking curve, while the middle panel presents the raw throttle signal and corresponding acceleration profile.
Because trajectory generation is performed state-wise—predicting only the immediate next action—the reference speed may vary discontinuously over time. These discontinuities are amplified by the PID controller, whose binary throttle/brake responses induce abrupt changes in acceleration, ultimately leading to discomfort during vehicle operation. Such execution-level instabilities are a major contributor to the discomfort observed across CBV methods.
While many CBV methods attempt to mitigate discomfort through fine-tuning strategies that incorporate post-action feedback via reward shaping or expert action alignment—RIFT adopts a different approach. It employs a state-wise reward model (see Section B.6) that quantifies comfort within the trajectory’s virtual time domain.
To further analyze this, we visualize both the actual acceleration after executing a selected trajectory and the corresponding virtual-time acceleration (shown in Figure 7). The results reveal that while virtual-time acceleration aligns with actual motion at the beginning of the trajectory, it underestimates acceleration variations in later segments. This leads to an overly conservative estimation of trajectory-level discomfort, resulting in insufficient supervision during training and reflected in RIFT’s comfort performance in Table 7.
In summary, the discomfort exhibited by CBV methods can be attributed to two primary sources:
  • Tracking instability, caused by discontinuities in planned trajectories and the limited control fidelity of PID controllers. Discrete, state-wise planning combined with low-resolution, often binary control outputs amplifies acceleration fluctuations and leads to uncomfortable motion.
  • Inadequate comfort modeling, particularly in state-wise reward formulations such as that adopted by RIFT and GPRO. These formulations fail to capture long-term trajectory-level discomfort, leading to insufficient supervision during training and suboptimal comfort performance.
D.3 Visualization of the AV-Centered Closed-Loop Simulation
To qualitatively evaluate the robustness of RIFT across diverse AV-centered scenarios, we provide additional temporal visualizations of closed-loop simulations. As shown in Figure 8, the traffic scene consists of the autonomous vehicle (AV, controlled by PDM-Lite), background vehicles (BVs), and critical background vehicles (CBVs), which interact dynamically over time.
The visualizations demonstrate the ability of RIFT to generate temporally coherent, realistic, and controllable trajectories across a variety of traffic situations. Even under complex and evolving closed-loop conditions, RIFT maintains stable multimodal behavior, highlighting its effectiveness in simulating realistic and controllable traffic flows around the AV.
Figure 8: Extra temporal visualization of RIFT in the AV-centered closed-loop simulation environment. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
Appendix E Social Impact
Positive Societal Impacts.
This work introduces a practical framework that bridges realism and controllability in traffic simulation. By decoupling the pre-training and fine-tuning stages, our method enables models pre-trained on real-world datasets to be effectively adapted to physics-based simulators, preserving trajectory-level realism while improving long-horizon, closed-loop performance. This paradigm offers a viable path for data-driven approaches to transition toward physics-based simulators, facilitating more reliable closed-loop testing and training. Consequently, it contributes to the development of safer and more robust autonomous systems.
Negative Societal Impacts.
While fine-tuning in physics-based simulators improves closed-loop performance, it may also lead to overfitting to the specific characteristics of the simulator. As a result, the learned policy could struggle to generalize beyond the simulated environment, giving rise to a sim-to-real gap. This gap poses challenges for real-world deployment, as models that perform well in simulation may not retain the same level of reliability when applied to actual autonomous driving systems. Such discrepancies can affect the testing and training stages, highlighting the need for further work to ensure real-world transferability.