### v3

# RIFT: Group-Relative RL Fine-Tuning for Realistic and Controllable Traffic Simulation

Abstract
Achieving both realism and controllability in closed-loop traffic simulation remains a key challenge in autonomous driving. Dataset-based methods reproduce realistic trajectories but suffer from covariate shift in closed-loop deployment, compounded by simplified dynamics models that further reduce reliability. Conversely, physics-based simulation methods enhance reliable and controllable closed-loop interactions but often lack expert demonstrations, compromising realism. To address these challenges, we introduce a dual-stage AV-centric simulation framework that conducts imitation learning pre-training in a data-driven simulator to capture trajectory-level realism and route-level controllability, followed by reinforcement learning fine-tuning in a physics-based simulator to enhance style-level controllability and mitigate covariate shift. In the fine-tuning stage, we propose RIFT, a novel group-relative RL fine-tuning strategy that evaluates all candidate modalities through group-relative formulation and employs a surrogate objective for stable optimization, enhancing style-level controllability and mitigating covariate shift while preserving the trajectory-level realism and route-level controllability inherited from IL pre-training. Extensive experiments demonstrate that RIFT improves realism and controllability in traffic simulation while simultaneously exposing the limitations of modern AV systems in closed-loop evaluation. Project Page: https://currychen77.github.io/RIFT/
1 Introduction
Reliable closed-loop traffic simulation is critical for developing advanced autonomous vehicle (AV) systems, supporting training and evaluation [1, 2]. An ideal traffic simulation should possess two key properties: realistic, reflecting real-world driving behavior; controllable, enabling customizable traffic simulation according to user requirements.
To balance these two essential properties, existing traffic simulation methods adopt different trade-offs depending on the underlying platform, often favoring either realism or controllability, as illustrated in Figure˜1. Methods based on data-driven simulators exploit real-world data to generate realistic trajectories by learning multimodal behavioral patterns through imitation learning (IL) [3, 4, 5, 6]. In addition to realism, recent studies on data-driven simulators have pursued controllability by conditioning scenario generation on user-specified inputs—such as text conditions [7, 8], goal conditions [9, 10], or cost functions [11, 12, 13]—producing scenarios that are both realistic and aligned with user requirements. However, their open-loop training paradigm introduces the covariate shift problem during closed-loop deployment, arising from the distribution mismatch between training and deployment states. Moreover, data-driven simulators often adopt simplified environment dynamics [14, 15], resulting in unrealistic interactions and state transitions that further degrade closed-loop reliability. In contrast, physics-based simulators provide fine-grained control over scenario configuration through physical engines, enabling high-fidelity closed-loop interactions. Nonetheless, the absence of expert demonstrations makes it challenging to reproduce realistic behavior. To mitigate this, several approaches employ reinforcement learning (RL) to directly acquire controllable behaviors through interaction with the simulator [16, 17, 18, 19], although often at the cost of realism. Other approaches enhance realism by injecting real-world traffic data into physics-based simulators [20, 21], but typically rely on log-replay or rule-based simulation, limiting controllability and interactivity. Despite recent advances, a fundamental trade-off persists between realism and controllability across both paradigms, making it challenging to achieve both simultaneously in interactive closed-loop scenarios.
Figure 1: Traffic Simulation across Different Platforms. (a) Data-driven Simulator: employs imitation learning to replicate real-world driving behaviors, but suffers from covariate shift and simplified dynamics; (b) Physics-based Simulator: enables controllable scenario construction via high-fidelity closed-loop interaction, but lacks large-scale real-world data; (c) Our framework: combines IL pre-training in a data-driven simulator to ensure realism with RL fine-tuning in a physics-based simulator to enhance controllability.
Drawing inspiration from the widely adopted “pre-training and fine-tuning” paradigm in large language models (LLMs) [22, 23, 24], we combine the strengths of two platforms. Specifically, we perform IL pre-training in a data-driven simulator to capture realism, followed by RL fine-tuning in a physics-based simulator to address covariate shift and enhance controllability.
Building on this insight, we propose a dual-stage AV-centric simulation framework (Figure˜1) that unifies the strengths of data-driven and physics-based simulators through a “pre-training and fine-tuning” paradigm, balancing realism and controllability in traffic simulation. In Stage 1, we pre-train a planning model via IL to generate realistic and multimodal trajectories conditioned on given route-level reference lines. This stage achieves both trajectory-level realism, capturing realistic and multimodal behavior patterns, and route-level controllability, guaranteeing compliance with prescribed reference lines. In Stage 2, we identify critical background vehicles (CBVs) through route-level interaction analysis, focusing on those most likely to interact with the AV. For these CBVs, we leverage the IL pre-trained model from Stage 1, conditioned on their route-level reference lines, to automatically generate realistic and multimodal trajectories that remain route-level controllable. On top of these generated candidates, we introduce RIFT, a novel group-relative RL fine-tuning strategy that improves controllability over driving styles and mitigates covariate shift. Unlike prior methods [25, 26] that fine-tune only the best trajectory or action, RIFT evaluates all candidate modalities via group-relative formulation [24] and employs a surrogate objective for stable optimization, enhancing style-level controllability and alleviating covariate shift while preserving the trajectory-level realism and route-level controllability established in Stage 1.
Our contributions can be summarized as:
  • We propose a dual-stage AV-centric simulation framework that combines IL pre-training in a data-driven simulator and RL fine-tuning in a physics-based simulator, leveraging their complementary strengths to balance realism and controllability.
  • We propose RIFT, a novel group-relative RL fine-tuning strategy that evaluates all candidate modalities through group-relative formulation and employs a surrogate objective for stable optimization, improving style-level controllability and alleviating covariate shift, while retaining the trajectory-level realism and route-level controllability inherited from IL pre-training.
  • Extensive experiments demonstrate that RIFT enhances the realism and controllability of traffic simulation, effectively exposing the limitations of modern AV systems under closed-loop settings.
2 Related Work
Realistic Traffic Simulation. A variety of generative architectures have been explored for realistic traffic simulation [27, 28, 29], including conditional variational autoencoders [30, 31, 32] and diffusion-based models [33, 34, 35, 36]. However, maintaining long-term stability remains challenging due to the covariate shift between open-loop training and closed-loop deployment. Recent methods such as SMART [37], GUMP [38], Trajeglish [39], and MotionLM [40] address this issue by formulating traffic simulation as a next-token prediction (NTP) task, leveraging discrete action spaces to improve closed-loop robustness. Despite these advances, most approaches remain confined to data-driven simulation platforms [14, 15, 41, 42], which typically adopt simplified environment dynamics. Such oversimplifications limit the reliability of long-term closed-loop interactions, especially in complex and interactive scenarios.
Controllable Traffic Simulation. Recent studies have introduced diverse conditioning mechanisms to generate traffic scenarios aligned with user preferences. CTG [11] and MotionDiffuser [12] employ diffusion models conditioned on cost-based signals. Language-conditioned methods, including CTG++ [13], LCTGen [8], and ProSim [9], enable user specification through language prompts. Other strategies adopt guided sampling (SceneControl [43]), retrieval-based generation (RealGen [44]), or reward-driven causality modeling (CCDiff [45]). Despite improving controllability, existing approaches remain confined to open-loop settings or simplified dynamics, and primarily target low-level control. High-level attributes such as driving style are underexplored, leaving the integration of realism and controllability in closed-loop simulation an open challenge.
Closed-Loop Fine-Tuning. Covariate shift—the mismatch between open-loop training and closed-loop deployment—remains a key challenge for reliable long-term traffic simulation. To address this, recent work explores fine-tuning strategies in the closed-loop setting. Hybrid IL and RL methods [25, 26, 46] enhance robustness but typically fine-tune the entire model via RL, which often compromises realism due to the difficulty of designing human-aligned reward functions. Supervised fine-tuning approaches such as CAT-K [47] show strong performance but rely on expert demonstrations, limiting scalability. TrafficRLHF [48] improves alignment through reinforcement learning with human feedback (RLHF), but demands costly human input and suffers from reward model instability. Moreover, most existing methods focus on optimizing the best action or trajectory, ignoring the inherent multimodality of traffic simulation, thus limiting behavioral diversity during fine-tuning.
3 Background
3.1 Task Redefinition
Following the widely adopted paradigm for closed-loop training and evaluation in autonomous driving [49, 50], our simulation framework includes a single autonomous vehicle (AV) navigating a predefined global route, accompanied by multiple rule-based background vehicles (BVs), forming an AV-centric closed-loop simulation environment. These BVs either provide diverse interactive data for training or serve to evaluate the AV’s robustness. Building upon this setup, we identify a subset of critical background vehicles (CBVs) that are more likely to interact with the AV. For these CBVs, the rule-based control is replaced with a well-trained planning model, enabling the synthesis of realistic and controllable behaviors in interactive closed-loop scenarios.
3.2 CBV-Centric Realistic Trajectory Generation
With recent advances in imitation learning, data-driven approaches have demonstrated strong performance in generating realistic, multimodal trajectories [51, 52, 53, 54, 55]. In fully observable simulation environments, Pluto [56] produces reliable, realistic, and multimodal trajectories by leveraging ground-truth states, while enabling route-level controllability through reference line encoding. These capabilities make Pluto a suitable choice for our planning model.
CBV-Centric Scene Encoding. Following [56], for each CBV in the scene, we extract its current feature $F_{\mathrm{cbv}}$, the historical features of neighboring vehicles $F_{\mathrm{neighbor}}$, and vectorized map features $F_{\mathrm{map}}$. These features are encoded into $E_{\mathrm{cbv}}\in\mathbb{R}^{1\times D}$, $E_{\mathrm{neighbor}}\in\mathbb{R}^{N_{\mathrm{neighbor}}\times D}$, and $E_{\mathrm{map}}\in\mathbb{R}^{N_{\mathrm{map}}\times D}$, respectively, where $N_{\mathrm{neighbor}}$ and $N_{\mathrm{map}}$ denote the number of neighboring vehicles and map elements, and $D$ is the embedding dimension. To model the interactions among these embeddings, we concatenate them and apply a global positional embedding ($\mathrm{PE}$) to obtain the unified scene embedding $E_{s}\in\mathbb{R}^{(1+N_{\mathrm{neighbor}}+N_{\mathrm{map}})\times D}$ as:
$E_{s}$ $=\mathrm{concat}(E_{\mathrm{cbv}},E_{\mathrm{neighbor}},E_{\mathrm{map}})+\mathrm{PE}.$ (1)
This scene embedding $E_{s}$ is then passed through $N$ Transformer encoder blocks for feature aggregation, yielding the final CBV-centric scene embedding $E_{\mathrm{enc}}$. Each encoder block follows the standard Transformer formulation. Specifically, the $i$-th block is defined as:
$E^{i}_{s}=E^{i-1}_{s}+\mathrm{MHA}(\mathrm{LayerNorm}({E^{i-1}_{s}})),$ (2)
$E^{i}_{s}=E^{i}_{s}+\mathrm{FFN}\left(\mathrm{LayerNorm}(E^{i}_{s})\right),$
where $\mathrm{MHA}$ is the standard multi-head attention function, $\mathrm{FFN}$ is the feedforward network layer.
Multimodal Trajectory Decoding. To capture the multimodal nature of real-world driving behaviors, we adopt the longitudinal-lateral decoupling mechanism proposed in [56]. This approach leverages reference line information to construct high-level lateral queries $Q_{\mathrm{lat}}\in\mathbb{R}^{N_{\mathrm{ref}}\times D}$, and introduces learnable longitudinal queries $Q_{\mathrm{lon}}\in\mathbb{R}^{N_{\mathrm{lon}}\times D}$. These are concatenated and projected to form the multimodal navigation query $Q_{\mathrm{nav}}\in\mathbb{R}^{N_{\mathrm{ref}}\times{N_{\mathrm{lon}}}\times D}$ as:
$Q_{\mathrm{nav}}=\mathrm{Projection}(\mathrm{concat}(Q_{\mathrm{lat}},Q_{\mathrm{lon}})),$ (3)
where $N_{\mathrm{ref}}$ and $N_{\mathrm{lon}}$ denote the number of reference lines and longitudinal anchors, respectively. The navigation query $Q_{\mathrm{nav}}$ and the scene embedding $E_{\mathrm{enc}}$ are then fed into $N$ decoder blocks to model lateral, longitudinal, and cross-modal interactions. Each decoder block is structured as:
$\hat{Q}^{i-1}_{\mathrm{nav}}$ $=\mathrm{SelfAttn}(\mathrm{SelfAttn}(Q^{i-1}_{\mathrm{nav}},\mathrm{dim}=0),\mathrm{dim}=1),$ (4)
$Q^{i}_{\mathrm{nav}}$ $=\mathrm{CrossAttn}(\hat{Q}^{i-1}_{\mathrm{nav}},E_{\mathrm{enc}},E_{\mathrm{enc}}).$
$\mathrm{SelfAttn}$, $\mathrm{CrossAttn}$ denote multi-head self-attention and cross-attention, respectively. Given the decoder’s final output $Q_{\mathrm{dec}}$, two MLP heads are applied to produce the CBV-centric multimodal trajectories $\mathcal{T}\in\mathbb{R}^{N_{\mathrm{ref}}\times N_{\mathrm{lon}}\times T\times 6}$ and their confidence scores $\mathcal{S}\in\mathbb{R}^{N_{\mathrm{ref}}\times N_{\mathrm{lon}}}$:
${\mathcal{T}}=\mathrm{MLP}(Q_{\mathrm{dec}}),\;{\mathcal{S}}=\mathrm{MLP}(Q_{\mathrm{dec}}),$ (5)
where $T$ is the prediction horizon, and each trajectory point $\tau_{t}^{i}$ encodes $p_{x},p_{y},\cos{\theta},\sin{\theta},v_{x},v_{y}]$.
4 Methodology
Leveraging the IL pre-trained planning model described in Section˜3.2, realistic and multimodal trajectories can be generated across diverse scenarios conditioned on reference lines. However, the open-loop training paradigm leaves the policy vulnerable to covariate shift, even with contrastive learning [57, 58] or data augmentation [56]. To address this, we propose RIFT, a group-relative RL fine-tuning strategy that enhances style-level controllability and mitigates covariate shift while preserving the trajectory-level realism and route-level controllability from pre-training. The following sections detail RIFT’s implementation within the physics-based simulator.
Figure 2: Overview of the RIFT: Building on the IL pre-trained model, RIFT performs route-level interaction analysis to identify critical background vehicles and the associated reference lines, enabling the generation of realistic and multimodal trajectories. To isolate style-level controllability from the trajectory-level realism and route-level controllability established during pre-training, only the scoring head is fine-tuned via RIFT while freezing other components. Specifically, RIFT computes group-relative advantages over all candidate rollouts, promoting alignment with user-preferred styles and mitigating covariate shift through RL fine-tuning.
4.1 Route-Level Interaction Analysis
Following [1], we address the “curse of rarity” [59] by selectively intervening in a set of critical background vehicles (CBVs) at key moments, while keeping non-critical agents under rule-based control for efficiency. CBVs are identified via route-level interaction analysis between the AV’s predefined global route and the candidate routes of surrounding vehicles, selecting the vehicle with the highest interaction probability (details in Section˜B.2).
The corresponding route-level reference line is then used as a condition for the IL pre-trained planning model (Section˜3.2) to synthesize realistic and multimodal trajectories. For each identified CBV, the model generates $N_{\mathrm{ref}}\times N_{\mathrm{lon}}$ candidate trajectories, from which the highest-scoring one is selected for closed-loop execution. This process promotes realistic route-level interactions with the AV and enables the construction of meaningful interactive scenarios.
4.2 Group-Relative RL Fine-Tuning
Open-loop IL pre-training offers trajectory-level realism and route-level controllability; however, it inevitably suffers from covariate shift in closed-loop deployment, causing error accumulation and unrealistic long-term behaviors. Existing RL [60] and hybrid IL–RL methods [26] partially mitigate covariate shift, but their optimization is restricted to the executed rollout, disregarding alternative candidates and degrading multimodality. More critically, covariate shift induces asymmetric degradation across model components: under the generation–selection paradigm, the generation head, conditioned on route-level priors, remains robust and consistently produces realistic multimodal candidates, whereas the scoring head, trained solely through imitation, is more vulnerable to distribution mismatch. These challenges motivate three key requirements for fine-tuning: (i) preserving multimodality, (ii) addressing asymmetric covariate shift, and (iii) ensuring stable policy improvement. We address these requirements through a unified framework that combines group-relative optimization, asymmetry-aware fine-tuning, and dual-clip stabilization.
To preserve multimodality, we adopt group-relative formulation [24], which evaluates all candidate modalities within the group and assigns higher relative advantages to those better aligned with user-preferred styles. Considering closed-loop dynamics, we evaluate simulated rollouts rather than raw trajectories to mitigate plan–rollout deviation. Specifically, given $\smash{G=N_{\mathrm{ref}}\times N_{\mathrm{lon}}}$ candidate trajectories $\smash{\mathcal{T}=\{\tau_{i}\}_{i=1}^{G}}$ for a CBV at state $s$, we conduct forward simulation [61] (see Appendix B.6) to obtain rollouts $\smash{\widetilde{\mathcal{T}}=\{\widetilde{\tau}_{i}\}_{i=1}^{G}}$. Each rollout is evaluated by a user-defined state-wise reward model $\mathrm{StateWiseRM}$, yielding the corresponding discounted returns $\smash{\mathcal{R}=\{R_{i}\}_{i=1}^{G}}$ from which we derive the group-relative advantages $\smash{\mathcal{A}=\{\hat{A}_{i}\}_{i=1}^{G}}$ as follows:
$R_{i}(s)=\sum_{t=0}^{T}\gamma^{t}\left[\mathrm{StateWiseRM}(\widetilde{\tau}_{i}^{t},s)\right],\ \qquad\hat{A}_{i}(s)=\frac{R_{i}(s)-\mathrm{mean}(\mathcal{R})}{\sqrt{\mathrm{Var}(\mathcal{R})+\varepsilon}}.$ (6)
Here, $\hat{A}_{i}$ quantifies the performance of each rollout relative to the group, promoting high-return rollouts without suppressing alternative modes.
In standard GRPO [24], sampling from the old policy implicitly induces old-policy weighting. Extending this to our enumerated setting involves averaging terms weighted by $\pi_{\theta_{\mathrm{old}}}$ in conjunction with the importance ratio $\smash{\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}\!\mid\!s)/\pi_{\theta_{\mathrm{old}}}(\tau_{i}\!\mid\!s)}$, which yields a low-variance estimate of the old-policy expectation over the enumerated support. The aggregated objective is:
$\begin{split}\mathcal{J}(\theta)=\mathbb{E}_{s\sim\mathcal{D}}\Bigg[\sum_{i=1}^{G}\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)\min\left[\rho_{i}(\theta)\hat{A}_{i},\mathrm{clip}\left(\rho_{i}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right]\Bigg]-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right],\end{split}$ (7)
where $\smash{\pi_{\mathrm{ref}}}$ denotes the IL pre-trained model. While exact over the enumerated support, this scheme overemphasizes frequent modes and under-represents rare but high-return ones, causing mode collapse and reduced diversity. To balance modality contributions, we adopt an equal-weight objective:
$\begin{split}\mathcal{J}(\theta)=\mathbb{E}_{s\sim\mathcal{D}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\min\left[\rho_{i}(\theta)\hat{A}_{i},\mathrm{clip}\left(\rho_{i}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right]\Bigg]-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right].\end{split}$ (8)
Under equal weighting, $\rho_{i}(\theta)$ regulates candidate updates rather than serving as a pure importance weight, removing old-policy bias and yielding balanced updates that preserve multimodality.
To address asymmetric covariate shift, we freeze the generation head to retain trajectory-level realism and fine-tune only the scoring head to enhance style-level controllability. In this setting, constraining the scoring head with the KL term to the IL pre-trained model would anchor learning to a biased reference, thereby hindering adaptation. We therefore remove the KL term, allowing the scoring head to adapt freely while leveraging the stable candidates provided by the frozen generation head.
Removing the KL term improves flexibility but raises stability concerns. Although the clipped-ratio mechanism in PPO constrains update magnitude, it proves insufficient in the group-relative setting. Specifically, when a rare trajectory under the old policy receives a higher probability from the current policy despite a negative advantage, the product $\smash{\rho_{i}(\theta)\hat{A}_{i}}$ can become disproportionately large and destabilize learning. To address this, we incorporate the dual-clip surrogate from Dual-Clip PPO [62, 63], which lower-bounds clipped negative advantages. This establishes a trust-region-like constraint that guarantees bounded per-candidate updates (see Lemma˜A.3), thereby preventing extreme negative shifts while preserving responsiveness to user-preferred styles. The resulting surrogate objective, termed RIFT, is
$\mathcal{J}_{\mathrm{RIFT}}(\theta)$ $=\mathbb{E}_{s\sim\mathcal{D}}\!\left[\frac{1}{G}\sum_{i=1}^{G}\;\psi\big(\rho_{i}(\theta),\,\hat{A}_{i}\big)\right],$ (9)
[…]
This objective integrates multimodality preservation, asymmetry-aware fine-tuning, and stable optimization into a unified framework, enhancing style-level controllability and mitigating covariate shift while retaining trajectory-level realism and route-level controllability (analysis in Appendix˜A).
5 Experiment
This section systematically addresses the following research questions: Q1: How does RIFT compare with representative baselines in terms of the realism and controllability of the generated traffic scenarios? Q2: How can the generated traffic scenario be effectively utilized to support downstream autonomous driving tasks? Q3: How do the components of RIFT contribute to overall performance, and to what extent is style-level controllability preserved under varying user-specified driving styles?
5.1 Experiment Setups
Under the dual-stage AV-centric simulation framework, we adopt Pluto [56] as our planning model for its well-established performance and open-source implementation. To ensure fair comparison, we use the official IL pre-trained checkpoint provided by Pluto, trained on the nuPlan dataset [15]. Simulations are conducted in CARLA [64], leveraging Bench2Drive [49] to support AV-centric closed-loop simulation and evaluation. Implementation details, training protocols, and evaluation settings are described in Appendix˜B.
Table 1: Comparison in Controllability and Realism. Metrics are evaluated under the PDM-Lite [65] AV setting across three random seeds, with the best and the second-best results highlighted accordingly.
Method Type Kinematic Metrics Interaction Metrics Map Metrics
S-SW $\uparrow$ S-WD $\downarrow$ A-SW $\uparrow$ CPK $\downarrow$ RP $\uparrow$ 2D-TTC $\uparrow$ ACT $\uparrow$ ORR $\downarrow$
Pluto IL 0.88 $\pm$ 0.01 5.81 $\pm$ 0.06 0.90 $\pm$ 0.01 5.06 $\pm$ 2.69 564.14 $\pm$ 114.41 2.50 $\pm$ 1.48 2.44 $\pm$ 1.39 0.24 $\pm$ 0.15
PPO RL 0.95 $\pm$ 0.01 4.45 $\pm$ 0.15 0.89 $\pm$ 0.02 13.95 $\pm$ 2.34 409.51 $\pm$ 30.38 2.59 $\pm$ 1.60 2.52 $\pm$ 1.57 9.17 $\pm$ 2.39
FREA RL 0.93 $\pm$ 0.01 5.10 $\pm$ 0.14 0.93 $\pm$ 0.01 30.42 $\pm$ 5.28 292.81 $\pm$ 68.54 2.71 $\pm$ 1.40 2.67 $\pm$ 1.41 9.01 $\pm$ 2.09
FPPO-RS RL 0.87 $\pm$ 0.01 5.80 $\pm$ 0.11 0.80 $\pm$ 0.03 21.39 $\pm$ 3.23 356.79 $\pm$ 26.19 2.55 $\pm$ 1.69 2.53 $\pm$ 1.68 8.60 $\pm$ 0.25
SFT-Pluto SFT 0.88 $\pm$ 0.02 6.01 $\pm$ 0.19 0.87 $\pm$ 0.02 6.33 $\pm$ 2.23 780.48 $\pm$ 41.05 2.20 $\pm$ 1.64 2.12 $\pm$ 1.51 0.06 $\pm$ 0.07
RS-Pluto SFT+RLFT 0.93 $\pm$ 0.00 5.40 $\pm$ 0.15 0.92 $\pm$ 0.01 4.11 $\pm$ 3.90 819.40 $\pm$ 74.07 2.27 $\pm$ 1.45 2.23 $\pm$ 1.43 1.05 $\pm$ 0.31
RTR-Pluto SFT+RLFT 0.85 $\pm$ 0.00 6.24 $\pm$ 0.16 0.81 $\pm$ 0.03 6.98 $\pm$ 2.59 481.60 $\pm$ 70.19 2.55 $\pm$ 1.60 2.47 $\pm$ 1.51 0.08 $\pm$ 0.09
PPO-Pluto RLFT 0.95 $\pm$ 0.01 4.96 $\pm$ 0.31 0.90 $\pm$ 0.02 6.89 $\pm$ 3.19 683.57 $\pm$ 38.12 2.66 $\pm$ 1.50 2.60 $\pm$ 1.43 0.07 $\pm$ 0.13
REINFORCE-Pluto RLFT 0.92 $\pm$ 0.01 5.63 $\pm$ 0.19 0.90 $\pm$ 0.02 6.98 $\pm$ 0.86 813.70 $\pm$ 24.76 2.39 $\pm$ 1.64 2.30 $\pm$ 1.55 1.37 $\pm$ 1.13
GRPO-Pluto RLFT 0.94 $\pm$ 0.04 4.96 $\pm$ 0.89 0.96 $\pm$ 0.00 7.24 $\pm$ 4.04 892.65 $\pm$ 65.27 2.65 $\pm$ 1.44 2.61 $\pm$ 1.48 0.10 $\pm$ 0.08
RIFT-Pluto (ours) RLFT 0.97 $\pm$ 0.01 4.46 $\pm$ 0.43 0.93 $\pm$ 0.01 6.83 $\pm$ 2.62 995.33 $\pm$ 84.62 2.74 $\pm$ 1.30 2.71 $\pm$ 1.32 0.36 $\pm$ 0.20
Figure 3: Speed and Acceleration Distribution. RL-based methods tend to be interactive but unnatural, whereas supervised methods are overly conservative. RIFT strikes a balance, yielding higher interactivity with realistic distributional profiles, reducing hesitation while maintaining safe interactions.
Figure 4: Temporal comparisons illustrating RIFT’s superior performance over other baselines under AV-centric closed-loop simulation. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
Baseline. To systematically evaluate the effectiveness of RIFT in traffic simulation, we compare it against the following baselines, with implementation details provided in Section˜B.5.
  • Pure RL/IL: Methods trained solely with RL or IL, without fine-tuning, including Pluto [56], as well as FREA, FPPO-RS, and PPO, all from [18].
  • RLFT/SFT: Methods that fine-tune the pre-trained Pluto model using either RL or supervised objectives, including PPO-Pluto [60], REINFORCE-Pluto [66], GRPO-Pluto [24], and SFT-Pluto.
  • Hybrid: Methods that combine RL and supervised fine-tuning, including RTR-Pluto [25] and RS-Pluto [26].
All methods are fine-tuned on the scoring head to ensure fair comparisons, while isolating style-level controllability from trajectory-level realism and route-level controllability, as confirmed by the ablation studies in Section˜5.4. Following the realism standards of the Sim Agent Challenge in WOSAC [42], we adopt a normal style reward for all RL-based baselines, with details in Section˜B.7. Results under an aggressive style reward are reported in Section˜5.4.
Metrics. Building on the WOSAC evaluation framework, we categorize our evaluation metrics into three groups: kinematic metrics, interaction metrics, and map metrics. Kinematic metrics capture distributional motion properties (S-SW, S-WD, A-SW), as in [67], with the absence of ground-truth trajectories in CARLA precluding displacement-based measures (e.g., ADE, FDE). Interaction metrics evaluate agent interactions through collision frequency (Collision Per Kilometer, CPK), driving efficiency (Route Progress, RP), and safety-critical measures, including 2D-TTC [68] and ACT [69]. Map metrics evaluate adherence to road geometry through the Off-Road Rate (ORR). Collectively, these metrics comprehensively evaluate realism and controllability in closed-loop simulation; detailed definitions are in Section˜B.8.
5.2 Realistic and Controllable Traffic Scenario Generation (Q1)
Main Results. To address Q1, we evaluate the controllability and realism of the generated scenario across CBV methods, with results summarized in Table˜1. RIFT consistently outperforms all baselines in both aspects across most settings. While supervised learning methods achieve slightly lower CPK and ORR, this improvement is primarily due to their inherently conservative behavior, derived from the expert PDM-Lite [65], which prioritizes safety by avoiding risky maneuvers.
This conservative tendency is further highlighted in Figure˜3, where supervised policies exhibit significantly lower speed and acceleration profiles. In contrast, RIFT strikes a more favorable balance between safety and interactivity. It achieves superior safety performance, as reflected by higher 2D-TTC and ACT scores, while avoiding the overly cautious behaviors typical of supervised approaches. As shown in Figure˜3, RIFT demonstrates higher average speed and acceleration, indicating more interactive behavior, while maintaining realistic motion profiles.
Qualitative Results. To further demonstrate the effectiveness of RIFT, we compare closed-loop simulations against representative baselines, as shown in Figure˜4. Baseline methods often suffer from unstable or low-quality trajectory selection in closed-loop settings, whereas RIFT consistently selects smooth, high-quality trajectories with superior temporal consistency. Further qualitative examples are presented in Section˜D.3.
5.3 Generated Traffic Scenarios for Closed-Loop AV Evaluation (Q2)
To address Q2, we assess the suitability of traffic scenarios generated by different CBV methods for closed-loop AV evaluation. Following KING [17], we adopt PDM-Lite [65]—a rule-based planner with privileged access—as a reference to evaluate two key scenario properties: feasibility, measured by Driving Score (DS), and naturalness, captured by our proposed Blocked Rate (BR). A high DS indicates that the AV can reliably complete the scenario, while a low BR reflects realistic interactions without excessive obstruction from surrounding vehicles. Together, DS and BR offer a principled basis for evaluating scenario quality.
To further assess the ability of each scenario to reveal weaknesses in learning-based planners, we compare PlanT [70], UniAD [53], and VAD [54] with PDM-Lite. As these models are sensitive to subtle or adversarial interactions, informative scenarios should induce noticeable performance drops. As shown in Table˜3, traffic generated by RIFT achieves the highest DS and lowest BR under PDM-Lite, while also causing the largest degradation across all learning-based planners. These results confirm that RIFT generates interactive and feasible scenarios that effectively expose limitations of modern AV systems. See Appendix˜C for detailed results.
Table 2: Comparison of AV Evaluation across CBV Methods. Each metric is evaluated across three random seeds, with the best and the second-best results highlighted accordingly.
Method PDM-Lite PlanT UniAD VAD
DS $\uparrow$ BR $\uparrow$ DS $\Delta{\text{DS}}\downarrow$ DS $\Delta{\text{DS}}\downarrow$ DS $\Delta{\text{DS}}\downarrow$
Pluto 77.84 $\pm$ 2.20 23.33 $\pm$ 5.77 42.52 $\pm$ 4.72 -35.32 73.73 $\pm$ 1.24 -4.11 66.87 $\pm$ 2.11 -10.97
PPO 76.26 $\pm$ 0.12 30.00 $\pm$ 0.00 36.39 $\pm$ 1.11 -39.87 69.79 $\pm$ 1.41 -6.47 67.64 $\pm$ 1.27 -8.62
FREA 83.53 $\pm$ 0.13 20.00 $\pm$ 0.00 39.61 $\pm$ 1.34 -43.92 69.29 $\pm$ 5.22 -14.24 67.57 $\pm$ 5.37 -15.96
FPPO-RS 83.52 $\pm$ 0.09 20.00 $\pm$ 0.00 38.85 $\pm$ 4.91 -44.67 75.13 $\pm$ 5.18 -8.39 69.15 $\pm$ 2.79 -14.37
SFT-Pluto 86.09 $\pm$ 2.04 13.33 $\pm$ 5.77 39.41 $\pm$ 4.97 -47.28 77.49 $\pm$ 5.93 -9.20 68.89 $\pm$ 0.87 -17.80
RS-Pluto 89.32 $\pm$ 1.41 13.33 $\pm$ 5.77 42.05 $\pm$ 4.08 -47.27 80.62 $\pm$ 0.78 -8.70 69.48 $\pm$ 5.02 -19.84
RTR-Pluto 87.64 $\pm$ 1.56 10.00 $\pm$ 0.00 40.08 $\pm$ 2.38 -47.56 77.69 $\pm$ 2.82 -9.95 66.27 $\pm$ 4.53 -21.37
PPO-Pluto 85.63 $\pm$ 2.02 16.67 $\pm$ 5.77 41.86 $\pm$ 2.78 -43.77 77.14 $\pm$ 3.36 -8.49 68.62 $\pm$ 3.16 -17.01
REINFORCE-Pluto 92.17 $\pm$ 3.45 10.00 $\pm$ 10.00 45.25 $\pm$ 1.75 -46.92 79.89 $\pm$ 1.97 -12.28 70.28 $\pm$ 3.58 -21.89
GRPO-Pluto 89.86 $\pm$ 2.10 6.67 $\pm$ 5.77 47.24 $\pm$ 5.67 -42.62 81.02 $\pm$ 0.64 -8.84 72.55 $\pm$ 0.74 -17.31
RIFT-Pluto (ours) 94.78 $\pm$ 1.37 0.00 $\pm$ 0.00 44.28 $\pm$ 3.15 -50.50 73.79 $\pm$ 6.53 -20.99 68.24 $\pm$ 3.23 -26.54
Table 3: Ablation Study on RIFT. Evaluation under PDM-Lite AV setting with three random seeds.
Method Kinematic Metrics Interaction Metrics Map Metrics
S-SW $\uparrow$ S-WD $\downarrow$ A-SW $\uparrow$ CPK $\downarrow$ RP $\uparrow$ 2D-TTC $\uparrow$ ACT $\uparrow$ ORR $\downarrow$
w/ Old-Weight 0.82 (-0.15) 6.24 (+1.78) 0.85 (-0.08) 7.51 (+0.68) 574.51 (-420.82) 2.70 (-0.04) 2.68 (-0.03) 0.00 (-0.36)
w/ All-Head 0.96 (-0.01) 4.70 (+0.24) 0.94 (+0.01) 7.84 (+1.01) 827.12 (-168.21) 2.83 (+0.09) 2.76 (+0.05) 0.43 (+0.07)
w/ KL 0.93 (-0.04) 5.33 (+0.87) 0.90 (-0.03) 7.05 (+0.22) 815.06 (-180.27) 2.76 (+0.02) 2.73 (+0.02) 0.38 (+0.02)
w/ PPO-Clip 0.91 (-0.06) 5.92 (+1.46) 0.94 (+0.01) 2.03 (-4.80) 655.39 (-339.94) 2.57 (-0.17) 2.54 (-0.17) 0.04 (-0.32)
w/ Aggressive 0.97 (+0.00) 3.89 (-0.57) 0.94 (+0.01) 8.41 (+1.58) 1053.76 (+58.43) 2.93 (+0.19) 2.88 (+0.17) 0.91 (+0.55)
RIFT-Pluto (ours) 0.97 4.46 0.93 6.83 995.33 2.74 2.71 0.36
5.4 Ablation Study (Q3)
Building on the design choices introduced in Section˜4.2, we systematically ablate five components of RIFT: weighting scheme (Old-Weight vs. Equal-Weight), fine-tuning module (Scoring Head vs. All Head), KL regularization (w/ KL vs. w/o KL), policy clipping (Dual-Clip vs. PPO-Clip), and style preference (Normal vs. Aggressive). All experiments share identical settings, and results are reported in Table˜3.
Equal-Weight vs. Old-Weight. Replacing old-policy weighting with equal weighting eliminates the likelihood bias toward frequent modes and enables balanced updates across all candidates. This leads to improved exploitation of high-return rollouts and better multimodality preservation.
Scoring Head vs. All Head. Freezing the generation head is crucial for retaining trajectory-level realism and route-level controllability. Fine-tuning all heads (w/ All Head) disrupts the pre-trained generation head and slightly degrades realism metrics, whereas fine-tuning only the scoring head achieves better controllability without compromising realism.
w/ KL vs. w/o KL. Anchoring the scoring head to the IL pre-trained reference via KL regularization (w/ KL) constrains adaptation to a biased reference under asymmetric covariate shift. Removing this term improves controllability while maintaining realism, confirming that free adaptation of the scoring head yields more effective policy improvement.
Dual-Clip vs. PPO-Clip. Replacing dual-clip with standard PPO clipping (w/ PPO-Clip) results in overly conservative behaviors and reduced efficiency, as extreme negative updates can dominate and suppress positive learning signals. Dual-clip bounds such updates while preserving responsiveness to high-return rollouts, producing more realistic and efficient behavior.
Normal vs. Aggressive. Adopting a more aggressive reward that emphasizes efficiency increases route progress but also raises collision and off-road rates, illustrating the efficiency–safety trade-off. These results demonstrate that RIFT supports flexible style shaping while maintaining stability and multimodality. Additional qualitative insights on controllability are provided in Section˜D.1.
6 Conclusion
In this work, we propose a dual-stage AV-centric simulation framework that conducts IL pre-training in a data-driven simulator to capture trajectory-level realism and route-level controllability, followed by RL fine-tuning in a physics-based simulator to address covariate shift and enhance style-level controllability. During fine-tuning, we introduce RIFT, a novel group-relative RL fine-tuning strategy that evaluates all candidate modalities using the group-relative formulation combined with a surrogate objective for optimization, thereby enhancing style-level controllability and mitigating covariate shift, while preserving the trajectory-level realism and route-level controllability established in IL pre-training. Extensive experiments demonstrate that RIFT generates scenarios with superior realism and controllability, effectively revealing the limitations of modern AV systems and further bridging the gap between traffic simulation and reliable closed-loop evaluation. Due to space limitations, more discussion on limitations and future direction can be found in Appendix E.2.
Appendix
Appendix A Theoretical Analysis
A.1 Setting
For each $s\!\sim\!\mathcal{D}$, a frozen trajectory generation head yields $\mathcal{C}(s)=\{\tau_{i}\}_{i=1}^{G}$. The trajectory score head defines $\pi_{\theta}(\tau_{i}\mid s)$ on $\mathcal{C}(s)$. Finite-horizon simulation provides returns
$R_{i}(s)=\sum_{t=0}^{T}\gamma^{t}\,\mathrm{StateWiseRM}(\widetilde{\tau}_{i}^{t},s).$ (10)
Uniform (within-group) moments:
$\mu_{\mathrm{uni}}(s)=\tfrac{1}{G}\sum_{j=1}^{G}R_{j}(s),\qquad\sigma^{2}_{\mathrm{uni}}(s)=\tfrac{1}{G}\sum_{j=1}^{G}(R_{j}(s)-\mu_{\mathrm{uni}}(s))^{2}.$ (11)
Uniform, centered advantages:
$\hat{A}_{i}(s)=\frac{R_{i}(s)-\mu_{\mathrm{uni}}(s)}{\sqrt{\sigma_{\mathrm{uni}}^{2}(s)+\varepsilon}},\quad\frac{1}{G}\sum_{i=1}^{G}\hat{A}_{i}(s)=0.$ (12)
Let $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}\mid s)/\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)$. Define the RIFT surrogate
$\mathcal{J}_{\mathrm{RIFT}}(\theta)=\mathbb{E}_{s}\!\left[\frac{1}{G}\sum_{i=1}^{G}\psi\big(\rho_{i}(\theta),\hat{A}_{i}(s)\big)\right],$ (13)
with dual-clip kernel
$\psi(\rho,\hat{A})=\begin{cases}\min\!\big(\rho\,\hat{A},~\mathrm{clip}(\rho,1-\epsilon,1+\epsilon)\,\hat{A}\big),&\hat{A}\geq 0,\\[3.0pt] \max\!\Big(\min\!\big(\rho\,\hat{A},~\mathrm{clip}(\rho,1-\epsilon,1+\epsilon)\,\hat{A}\big),~c\,\hat{A}\Big),&\hat{A}<0,\end{cases}\quad(\epsilon>0,~c>1).$ (14)
Assumptions.
(A) Support floor: there exists $\pi_{\min}>0$ such that $\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)\geq\pi_{\min}$ for all $(s,i)$, and $\pi_{\theta}>0\Rightarrow\pi_{\theta_{\mathrm{old}}}>0$ on $\mathcal{C}(s)$. (B) Boundedness: $|\hat{A}_{i}(s)|\leq A_{\max}$. (C) Regularity: $\log\pi_{\theta}(\tau_{i}\mid s)$ is $L$-Lipschitz and $C^{2}$ on compact $\Theta$.
A.2 Listwise View and Diversity Pressure
Consider the unclipped uniform surrogate
$L_{\mathrm{RIFT}}(\theta)=\mathbb{E}_{s}\!\left[\frac{1}{G}\sum_{i=1}^{G}\rho_{i}(\theta)\,\hat{A}_{i}(s)\right]=\mathbb{E}_{s}\!\left[\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{A}_{i}(s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)}\,\pi_{\theta}(\tau_{i}\mid s)\right].$ (15)
Proposition A.1 (Pairwise ascent and diversity).
Fix $s$ and shift an infinitesimal mass $\delta$ from $j$ to $i$ in $\pi_{\theta}(\cdot\mid s)$. Then $\delta L_{\mathrm{RIFT}}(\theta)=\tfrac{\delta}{G}\!\left(\tfrac{\hat{A}_{i}(s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)}-\tfrac{\hat{A}_{j}(s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{j}\mid s)}\right)$. Hence ascent moves mass toward larger $\hat{A}/\pi_{\mathrm{old}}$, amplifying underrepresented high-quality candidates when $\pi_{\mathrm{old}}$ is peaky.
Corollary A.2 (Top-1 Fisher consistency under uniform reference).
If $\pi_{\theta_{\mathrm{old}}}$ is uniform on $\mathcal{C}(s)$ and $i^{\star}(s)=\arg\max_{i}\hat{A}_{i}(s)$ is unique, any global maximizer of $L_{\mathrm{RIFT}}$ concentrates $\pi_{\theta}(\cdot\mid s)$ on $i^{\star}(s)$.
A.3 Clipping as Stability Control
Clipping is a pointwise pessimistic transform: for any $x=\rho\,\hat{A}$,
$\min(\rho\,\hat{A},\mathrm{clip}(\rho)\hat{A})\leq\rho\,\hat{A}\$.
Summed over mixed signs, there is no global monotone lower bound for $L_{\mathrm{RIFT}}$; instead, clipping serves to bound the value and the gradient.
Lemma A.3 (Bounded values and gradients).
If $|\hat{A}|\leq A_{\max}$, then for all $(s,i)$: (i) Value bounds: $\psi\!\in[0,(1+\epsilon)\hat{A}$] for $\hat{A}\!\geq\!0$, and $\psi\!\in[c\,\hat{A},0]$ for $\hat{A}\!<\!0$. (ii) Gradient bounds:
$\bigg|\frac{\partial\psi}{\partial\log\pi_{\theta}}\bigg|\ \leq\ \begin{cases}(1+\epsilon)|\hat{A}|,&\hat{A}\geq 0,\\ c\,|\hat{A}|,&\hat{A}<0,\end{cases}$
and on the negative branch when the dual-clip is active ($\psi=c\,\hat{A}$) the derivative is 0.
A.4 Smoothness w.r.t. Policy Divergence
Write $w_{i}(s)=\hat{A}_{i}(s)/\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)$ and note $|w_{i}|\leq A_{\max}/\pi_{\min}$ under Assumption A with $\pi_{\min}=\inf_{s,i}\pi_{\theta_{\mathrm{old}}}(i\mid s)>0$ (label-smoothing in practice). Then the unclipped surrogate is linear in $\pi_{\theta}$:
$L_{\mathrm{RIFT}}(\theta)-L_{\mathrm{RIFT}}(\theta^{\prime})=\mathbb{E}_{s}\!\left[\frac{1}{G}\sum_{i}w_{i}(s)\big(\pi_{\theta}(i\mid s)-\pi_{\theta^{\prime}}(i\mid s)\big)\right].$
Lemma A.4 (Lipschitz continuity via KL).
For any $\theta,\theta^{\prime}$,
$\big|L_{\mathrm{RIFT}}(\theta)-L_{\mathrm{RIFT}}(\theta^{\prime})\big|\ \leq\ \frac{A_{\max}}{\pi_{\min}}\ \sqrt{2\,\mathbb{E}_{s}\!\big[\mathrm{KL}\big(\pi_{\theta}(\cdot\mid s)\,\|\,\pi_{\theta^{\prime}}(\cdot\mid s)\big)\big]}\,.$
Proof.
By Hölder and Pinsker: $|\sum_{i}w_{i}\Delta\pi|\leq\|w\|_{\infty}\,\|\Delta\pi\|_{1}\leq(A_{\max}/\pi_{\min})\sqrt{2\,\mathrm{KL}(\pi_{\theta}\|\pi_{\theta^{\prime}})}$, then average over $s$. ∎
Lemma A.5 (Lipschitz continuity of clipped surrogate).
Because $\partial\psi/\partial\pi_{\theta}(i|s)$ is bounded by $A_{\max}/\pi_{\min}$ whenever the active branch is differentiable,
$|\mathcal{J}_{\mathrm{RIFT}}(\theta)-\mathcal{J}_{\mathrm{RIFT}}(\theta^{\prime})|\ \leq\ \frac{A_{\max}}{\pi_{\min}}\ \sqrt{2\,\mathbb{E}_{s}\!\big[\mathrm{KL}\big(\pi_{\theta}(\cdot\mid s)\,\|\,\pi_{\theta^{\prime}}(\cdot\mid s)\big)\big]}.$
A.5 Variance and Enumeration
Let $f_{i}(s;\theta)=\psi(\rho_{i}(\theta),\hat{A}_{i}(s))$. Exact enumeration yields $\operatorname{Var}\!\big(\tfrac{1}{G}\sum_{i}f_{i}\,\big|\,s\big)=0$ (assuming $\widetilde{\tau}_{i}$ and their evaluations are fixed during the update; otherwise, environment randomness still induces nonzero variance), while sampling $i$ i.i.d. within the group gives conditional variance $\operatorname{Var}(f_{i}\mid s)/N$ for $N$ samples.
A.6 Convergence of Stochastic Ascent
Theorem A.6 (Convergence to a stationary point).
Under Assumptions A–C, with step sizes $\eta_{k}>0$, $\sum_{k}\eta_{k}=\infty$, $\sum_{k}\eta_{k}^{2}<\infty$, and unbiased bounded-variance stochastic subgradients, the iterates of stochastic subgradient ascent on $\mathcal{J}_{\mathrm{RIFT}}$ satisfy
$\liminf_{k\to\infty}\mathbb{E}\!\big[\operatorname{dist}\big(0,\partial^{C}\mathcal{J}_{\mathrm{RIFT}}(\theta_{k})\big)\big]=0,$
where $\partial^{C}$ denotes the Clarke generalized gradient.
Sketch.
By Lemma A.3, generalized gradients are uniformly bounded; regularity of $\log\pi_{\theta}$ on compact $\Theta$ implies Lipschitz continuity. Robbins–Monro / Kushner–Yin results for non-smooth stochastic approximation apply. ∎
A.7 Why RIFT Preserves Multimodality
By Proposition A.1, ascent compares $\hat{A}/\pi_{\mathrm{old}}$: under peaky $\pi_{\mathrm{old}}$, underrepresented high-$\hat{A}$ candidates receive stronger positive updates, preserving and enhancing diversity. In the special case $\pi_{\mathrm{old}}$ is (approximately) uniform, RIFTreduces to a listwise ranking ascent that directly promotes larger $\hat{A}$.
Appendix B Experimental Details
B.1 Experiment Framework
Our framework for reliable AV-centric closed-loop simulation is developed upon well-established traffic simulation platforms, notably the CARLA Leaderboard [71] and Bench2Drive [49], which serve as standard benchmarks in autonomous driving research. Traditionally, these platforms use predefined scenarios along the AV’s global route to evaluate the multi-dimensional performance of AV methods. In contrast, we replace these static scenarios with dynamically generated traffic flows by randomly spawning background vehicles around the AV’s global path and simulating their behavior using rule-based driving policies, as described in Section˜3.1. Through the CBV identification mechanism outlined in Section˜B.2, we naturally introduce interactions between the AV and CBVs, thereby generating continuous, interactive scenarios over time. This framework serves as the foundation for both the training and evaluation processes in this paper.
B.2 Route-level Analysis for CBV Identification
Identifying Critical Background Vehicles (CBVs) is essential to our AV-centric closed-loop simulation. Let $\mathcal{V}_{\text{AV}}$ denote the autonomous vehicle (AV), and $\mathcal{V}_{\text{BV}}=\{\mathcal{V}_{i}\}_{i=1}^{N}$ represent the set of background vehicles in the environment. The AV navigates along a predefined global route $\mathcal{P}=\{p_{k}\}_{k=1}^{M}$, where each $p_{k}$ corresponds to a waypoint along the route. The goal of CBV identification is to select background vehicles that are likely to share the AV’s destination and have similar estimated travel distance, thereby facilitating route-level interactions between the AV and CBVs. The primary criterion for identifying CBVs is the relative distance-to-goal difference between the AV and each background vehicle. This is mathematically expressed as:
$\left|\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{i})-\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{\text{AV}})\right|<\delta,$ (16)
where, $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{i})$ and $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{\text{AV}})$ denote the estimated travel distance required for the background vehicle $\mathcal{V}_{i}$ and the AV to reach waypoint $p_{k}$, respectively. The distance-to-goal for each vehicle is computed by determining the distance from its current position to the target waypoint $p_{k}$ using the A* global path planning algorithm [72]. A threshold $\delta$ is introduced to define the maximum allowable difference in distance-to-goal. A background vehicle is considered critical and included in the CBV set $\mathcal{C}$ if the absolute distance-to-goal difference between it and the AV is smaller than $\delta$.
This approach selects background vehicles whose destinations and estimated travel distances are sufficiently aligned with those of the AV, thereby ensuring meaningful and realistic route-level interactions. Once a CBV is identified, the planning path previously generated via A* during distance-to-goal estimation is directly adopted as its global navigation path, which is further transformed into the reference line for downstream CBV planning, naturally introducing route-level interactions between the AV and CBVs. The threshold $\delta$ serves as a tunable parameter to adjust the sensitivity of the CBV selection process. In this study, we set $\delta$ to $15m$ to achieve a balanced trade-off between sensitivity and selection accuracy.
B.3 Algorithm Framework
For clarity, we summarize the procedure of RIFT within our AV-centric closed-loop simulation framework in Algorithm˜1. The planning model is initialized from the IL pre-trained checkpoint provided by Pluto official codebase [^1] https://github.com/jchengai/pluto, followed by RL fine-tuning within the CARLA simulator [64] to generate realistic and controllable traffic scenarios.
Algorithm 1 Procedure for RIFT in the AV-Centric Closed-Loop Simulation Framework.
1:Input: IL pre-trained planning model $\pi_{\theta_{\text{init}}}$, buffer $\mathcal{D}$ $\triangleright$ IL pre-training (nuPlan [15])
2:planning model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$
3:for $\text{iteraction}=1,\dots,I$ do $\triangleright$ RL fine-tuning (CARLA [64])
4:  Update the old planning model $\pi_{\theta_{old}}\leftarrow\pi_{\theta}$
5:  while $\mathcal{D}$ not full do $\triangleright$ Collect rollout data
6:   for $\text{step}=1,\dots,T$ do
7:
Obtain $G$ candidate trajectories $\{\tau_{i}\}_{i=1}^{G}$ from $\pi_{\theta_{old}}$ for each CBV $\triangleright$ Policy inference
8:
Compute simulated rollouts $\{\widetilde{\tau}_{i}\}_{i=1}^{G}$ from $\{\tau_{i}\}_{i=1}^{G}$ $\triangleright$ Forward simulation
9:
Compute reward $\{R_{i}\}_{i=1}^{G}$, advantage $\{\hat{A}_{i}\}_{i=1}^{G}$ for each $\widetilde{\tau}_{i}$ with Equation˜6
10:
Store transition into buffer $\mathcal{D}$
11:   end for
12:  end while
13:  for RIFT $\text{iteration}=1,\dots,\mu$ do $\triangleright$ Policy fine-tuning
14:
Sample mini-batches transition from the buffer $\mathcal{D}$
15:
Update model $\pi_{\theta}$ by maximizing the RIFT objective (Equation˜9)
16:  end for
17:end for
18:Output: RL fine-tuned planning model
B.4 Training Details
We perform RL fine-tuning on selected modules of the IL pre-trained planning model (Pluto). As shown in the ablation results (Section˜5.4), fine-tuning only the trajectory scoring head achieves the best trade-off between realism and controllability. Accordingly, all fine-tuning baselines adopt this setting to ensure consistency and fair comparison. Our training framework is built on the open-source Lightning platform [^2] https://github.com/Lightning-AI/pytorch-lightning. Fine-tuning is conducted on $2\times$ Bench2Drive220, while evaluation is performed on dev10, both from the Bench2Drive project. All experiments are conducted on NVIDIA GeForce RTX 4090D GPUs, with each fine-tuning run taking approximately 8 hours on a single GPU. Detailed training setups and hyperparameter configurations are provided in Table˜6 and Table˜6.
B.5 Baselines Detailed Description
To comprehensively evaluate RIFT in an AV-centric closed-loop simulation environment, we compare it against a range of baselines, including pure imitation learning (IL), pure reinforcement learning (RL), and various fine-tuning approaches based on IL, RL, or their combination. We initialize all fine-tuning methods from the pre-trained Pluto checkpoint and fine-tune only the trajectory scoring head to preserve trajectory-level realism. The details of each baseline are summarized below.
  • Pluto [56] is an open-source IL-based planning framework for autonomous driving. It processes vectorized scene representations as input and outputs multimodal trajectories for downstream planning. In the AV-centric closed-loop simulation, the method directly uses a pre-trained checkpoint without additional fine-tuning.
  • FREA [18] is an RL-based approach designed to generate safety-critical yet AV-feasible scenarios. It incorporates a feasibility-aware training objective. In the AV-centric closed-loop simulation, FREA selects potential collision points along the AV’s global route as adversarial goals.
  • PPO [18] is a variant of FREA that focuses solely on generating safety-critical scenarios. Unlike FREA, it disregards the feasibility constraints of AV and treats adversariality as the only optimization objective.
  • FPPO-RS [18] is another FREA variant that integrates AV’s feasibility constraints into the reward shaping process, thereby balancing adversariality with scenario reasonability.
  • PPO-Pluto fine-tunes the pre-trained planning model using the PPO algorithm [60]. The fine-tuning follows the same reward structure as detailed in Section˜B.7, aligning with RIFT.
  • REINFORCE-Pluto employs the REINFORCE algorithm [66] to fine-tune the pre-trained Pluto model under the same reward design as detailed in Section˜B.7.
  • GRPO-Pluto utilizes the basic GRPO algorithm [24] for fine-tuning, employing the pre-trained Pluto model as the reference for KL regularization, while incorporating the standard PPO-Clip.
  • SFT-Pluto is a purely supervised fine-tuning approach, where PDM-Lite [65] serves as the expert model, providing supervision at the target speed level.
  • RTR-Pluto [25] is a hybrid framework combining imitation and reinforcement learning. While the original RTR utilizes human driving trajectories as supervision, our setting replaces this with PDM-Lite due to the lack of human-level demonstrations. The RL component uses sparse infraction-based rewards, consistent with the original RTR, and applies PPO for optimization.
  • RS-Pluto [26] also adopts a hybrid IL+RL paradigm, originally trained via REINFORCE using ground-truth supervision and sparse rewards to ensure safety and realism. In our adaptation, PDM-Lite substitutes the ground-truth expert, while the rest of the methodology remains unchanged.
B.6 Forward Simulation
Trajectory-based imitation learning often overlooks underlying system dynamics, leading to discrepancies between planned and executed behavior [73]. To address this issue, we perform a forward simulation for each candidate trajectory ${\tau}_{i}$ of the CBV, yielding a rollout $\widetilde{\tau}_{i}$. The simulation couples a PID controller for trajectory tracking with a kinematic bicycle model for state propagation. The PID controller is identical to that used during closed-loop execution, ensuring behavioral consistency between training and deployment. By evaluating rollouts rather than raw trajectories, we reduce this dynamics gap and obtain more reliable assessments.
In parallel, we also forecast the motions of surrounding actors. During data collection, the current actions $a^{\text{bg}}$ of surrounding actors are recorded. Following the rule-based forecasting scheme in [65], these actions are assumed constant over the forecast horizon and are used to advance surrounding states. The resulting actor forecasts are combined with the CBV rollouts to compute rewards, thereby ensuring that interaction effects with the environment are faithfully captured in evaluation.
While subsequent rollout is open-loop, the first transition is closed-loop. This step integrates (i) the same PID policy as in real execution, (ii) the observed current actions of surrounding actors, and (iii) a kinematic bicycle model that approximates CARLA’s single-step dynamics. Accordingly, the transition from $(s,a,a^{\text{bg}})$ to $s^{\prime}$ produces a reward consistent with the standard RL structure $(s,a)\to s^{\prime}\to r$. Subsequent rollout steps serve as open-loop estimates of longer-horizon outcomes, enriching evaluation while preserving closed-loop fidelity at the transition boundary.
B.7 State-Wise Reward Model Setup
To capture diverse human driving styles, we decompose driving behaviors into distinct reward components, following [74]. Different styles are constructed by combining weights assigned to each reward component (detailed in Table˜6), enabling a range of behaviors from aggressive to conservative. The total driving reward is defined as:
$R=R_{\text{collision}}+R_{\text{off-road}}+R_{\text{comfort}}+R_{\text{lane}}+R_{\text{velocity}}+R_{\text{timestep}}.$ (17)
The individual terms are described as follows:
  • $R_{\text{collision}}=-\left(\alpha_{\text{collision}}+\lvert v\rvert\right)\mathbbm{1}_{\mathrm{collision}}$: penalizes collisions, with higher penalties at higher speeds.
  • $R_{\text{off-road}}=-\alpha_{\text{boundary}}\mathbbm{1}_{\text{boundary}}$ $\alpha_{\text{boundary}}$: penalizes deviations from the drivable area.
  • $R_{\text{comfort}}=-\alpha_{\text{comfort}}\left(\mathbbm{1}_{\lvert a\rvert>4}+\mathbbm{1}_{\lvert\omega\rvert>4}\right)$ penalizes excessive acceleration and angular acceleration.
  • $R_{\text{l-align}}=\alpha_{\text{l-align}}\left(\min\left(\cos\left(\theta_{f}\right),0\right)+\alpha_{\text{vel-align}}\min\left(\cos\left(\theta_{f}\right)*v,0\right)+0.25\left(1-\frac{\lvert\theta_{f}\rvert}{\pi/2}\right)\right)$: guides the agent to follow the correct driving direction and remain parallel to the lane markings.
  • $R_{\text{l-center}}=-\alpha_{\text{l-center}}\left(\mathbbm{1}_{\cos\left(\theta_{f}\right)>0.5}*\left(\lvert x_{f}-\alpha_{\text{center-bias}}\rvert-\frac{0.05}{\exp\left(\lvert x_{f}-\alpha_{\text{center-bias}}\rvert-0.5\right)}\right)\right)$: guides the agent to prefer trajectories that remain centered within the lane.
  • $R_{\text{velocity}}=\alpha_{\text{velocity}}\max\left(\cos\left(\theta_{f}\right),0.0\right)\,\mathbbm{1}_{3<\lvert v\rvert<20}*\lvert v\rvert$: promotes forward movement and biases the agent toward choosing routes with consistent traffic flow rather than traffic jams.
  • $R_{\text{timestep}}=-\alpha_{\text{timestep}}\,\mathbbm{1}_{\lvert v\rvert>0\,\lor\,\lvert a\rvert>0}$ applies a small per-step penalty, encouraging efficiency. It is disabled when the agent is stationary to allow appropriate waiting behavior at intersections.
Building on the reward definitions above, we construct a state-wise reward model $\mathrm{StateWiseRM}(\cdot)$, which computes a scalar reward based on a set of interpretable features extracted from each rollout point $\widetilde{\tau}^{t}_{i}$. Specifically, we define a feature extraction function $\phi(\widetilde{\tau}^{t}_{i})$ as:
$\phi(\widetilde{\tau}^{t}_{i})=(\mathbbm{1}_{\text{collision}},\mathbbm{1}_{\text{boundary}},a_{\text{long}},a_{\text{lat}},\theta_{f},x_{f},v,a),$ (18)
where:
  • $\mathbbm{1}_{\text{collision}}$ and $\mathbbm{1}_{\text{boundary}}$ are binary indicators of potential collisions and off-road violations;
  • $a_{\text{long}}$ and $a_{\text{lat}}$ denote the longitudinal and lateral acceleration;
  • $v$ and $a$ are the magnitudes of velocity and acceleration;
  • $x_{f}$ is the lateral distance to the nearest lane centerline;
  • $\theta_{f}$ is the heading deviation concerning the lane direction.
The state-wise reward is then computed as:
$r^{t}_{i}=\mathrm{StateWiseRM}(\phi(\widetilde{\tau}^{t}_{i}),s).$ (19)
All features, except the infraction indicators, are directly derived from the rollouts. To estimate future infractions, we follow the forecasting model in [65] to simulate other agents’ future positions based on current states and actions and identify collisions via bounding box overlap. Off-road violations are detected by projecting the rollout trajectory onto the HD map and checking its occupancy relative to the drivable area polygon set.
Table 4: Hyperparameters used in RIFT Training.
Parameter Value
Batch size 256
Rollout buffer capacity 4096
Fine-tune initial LR $1\times e^{-4}$
Minimum LR $1\times e^{-6}$
LR decay across iteration 0.9
LR schedule Cosine
Num. RIFT epoch 16
Warmup Epoch of RIFT 3
AdamW weight-decay $1\times e^{-5}$
Table 5: Hyperparameters of RIFT or RL baselines.
Parameter Value
PPO clipping ratio $\epsilon$ 0.2
Dual-clip ratio $c$ 3
Discount factor $\gamma$ 0.98
$\lambda_{\text{GAE}}$ [75] 0.98
Hidden dimension $D$ 128
Num. lon. queries $N_{lon}$ 12
Traj. time horizon $T$ 80
Map radius 120m
Frame rate 10Hz
Table 6: Reward Parameters for Different Driving Styles.
Parameter Normal Aggressive
$\alpha_{\text{collision}}$ 20.0 5.0
$\alpha_{\text{boundary}}$ 5.0 5.0
$\alpha_{\text{comfort}}$ 0.8 0.8
$\alpha_{\text{l-align}}$ 0.5 0.5
$\alpha_{\text{vel-align}}$ 0.05 0.05
$\alpha_{\text{l-center}}$ 0.6 0.6
$\alpha_{\text{center-bias}}$ 0.0 0.0
$\alpha_{\text{velocity}}$ 0.1 0.2
$\alpha_{\text{timestep}}$ 0.1 0.1
B.8 Controllability and Realism Metrics
Kinematic Metrics. Following [42], kinematic realism is typically evaluated against ground-truth trajectories. As CARLA provides no expert demonstrations, we adopt distribution-level metrics [67, 76] to assess CBV behavior in terms of speed and acceleration. Specifically, we employ three measures—the Shapiro–Wilk test on speed (S-SW), the Wasserstein Distance on speed (S-WD), and the Shapiro–Wilk test on acceleration (A-SW)—defined as follows:
  • Wasserstein Distance (WD) [77]: measures the distance between two distributions $\mu$ and $\nu$. Since CARLA provides a predefined target speed for agents, we use WD to compare the simulated CBV speed distribution with the target speed distribution as the reference.
    $\text{WD}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\mathbb{E}_{(x,y)\sim\gamma}\left[d(x,y)\right].$ (20)
  • Shapiro–Wilk test (SW) [78]: evaluates the normality of speed and acceleration distributions—a simplifying assumption supported by empirical traffic studies [25, 79]—to capture the statistical naturalness of CBV motion.
    $\text{SW}=\frac{\left(\sum_{i=1}^{n}a_{i}x_{(i)}\right)^{2}}{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}},$ (21)
    where $a_{i}$ are coefficients, $x_{(i)}$ are the ordered data points, $x_{i}$ are the sample values, $\bar{x}$ is the sample mean, and $n$ is the number of data points.
Interaction Metrics. Following the metric design principles proposed in the WOSAC challenge [42] and other widely adopted evaluation frameworks [45, 67], we adopt a set of well-established metrics to comprehensively evaluate agent interactions:
  • Collision Per Kilometer (CPK) [67]: the average number of scenario collisions per kilometer of driving distance.
  • Route Progress (RP) [67]: the total distance traveled by all CBVs, reflecting route completion.
  • 2D Time-to-Collision (2D-TTC) [68]: the minimum of longitudinal and lateral time-to-collision from the AV’s perspective, capturing the interaction risk posed by CBVs.
  • Anticipated Collision Time (ACT) [69] : a safety-critical metric measuring the AV’s proximity to potential collisions, reflecting the interaction intensity introduced by CBVs.
Map Metrics. Map metrics evaluate adherence to road geometry, reflecting how well agents remain within drivable areas and comply with map constraints.
  • Off-Road Rate (ORR) [67]: the percentage of time CBVs spend off-road on average.
Appendix C AV Evaluation Details
C.1 AV Methods Implementation
To assess the effectiveness of RIFT in generating reliable and interactive scenarios for AV evaluation in the AV-centric closed-loop simulation environment, we evaluate the following representative and stable AV methods:
  • PDM-Lite [65]: A rule-based privileged expert method that achieves state-of-the-art performance on the CARLA Leaderboard 2.0 by leveraging components such as the Intelligent Driver Model and the kinematic bicycle model. This open-source method serves as a strong baseline for comparison.
  • PlanT [70]: An explainable, learning-based planning method that operates on an object-level input representation and is trained through imitation learning.
  • UniAD [53]: A planning‑oriented unified framework integrating perception, prediction, mapping, and planning into one end‑to‑end model using query-based interfaces.
  • VAD [54]: A fast, end‑to‑end vectorized driving paradigm representing scenes with vectorized motion and map elements for efficient, safe planning.
C.2 AV Evaluation Metrics
As detailed in Sections˜B.1 and B.4, we develop an AV-centric closed-loop simulation environment, including a training and evaluation pipeline based on Bench2Drive. The AV closed-loop evaluation metrics proposed in Bench2Drive extend the original metrics of the CARLA Leaderboard by emphasizing the specific strengths and weaknesses of different methods across various aspects, such as merging and overtaking, thereby making them suitable for evaluating performance under predefined scenarios. However, as noted in Section˜B.1, replacing predefined scenarios with CBV-generated traffic scenarios precludes the evaluation of specific AV capabilities. To systematically assess the quality of traffic scenarios generated by different CBV methods, we follow the practice of KING [17] and introduce PDM-Lite [65]—a rule-based privileged planner—as a reference AV. By measuring its performance under various CBV methods, we evaluate:
  • Feasibility, via PDM-Lite’s Driving Score (DS)—a high DS indicates the PDM-Lite can complete its route without severe collisions or rule violations, implying the generated traffic scenario is feasible.
  • Naturalness, via a newly proposed metric, Blocked Rate (BR)—a low BR suggests that CBVs do not unrealistically obstruct the AV, reflecting naturalistic behavior.
These metrics enable a principled comparison of traffic quality generated by different CBV methods. Furthermore, to assess the capacity of generated traffic scenarios to expose AV limitations, we test multiple learning-based AV methods under an identical CBV method and quantify their relative performance drop compared to PDM-Lite [65]. The relative driving score degradation ($\Delta{DS}$) reflects how effectively the traffic scenario stresses the AV policy, with larger drops indicating stronger capability in revealing planning weaknesses.
The evaluation metrics are summarized as follows:
  • Driving Score (DS): $R_{i}P_{i}$ — The main metric of the leaderboard, calculated as the product of route completion and the infraction penalty. Here, $R_{i}$ represents the percentage of completion of the $i$-th route, and $P_{i}$ denotes the infraction penalty. The maximum value is 100.
  • Block Rate (BR): The average number of occurrences where a CBV fails to navigate its route normally and obstructs the AV’s progress.
  • Relative Driving Score Degradation ($\Delta{DS}$): The reduction in Driving Score of a learning-based AV compared to PDM-Lite under the same CBV method, indicating how effectively the scenario reveals weaknesses in AV planning.
C.3 End-to-End AV Visualization
To further validate the feasibility of RIFT as a closed-loop evaluation framework, we extend its application beyond traffic scenario generation to testing end-to-end autonomous driving algorithms. In contrast to conventional adversarial approaches that often introduce unrealistic or overly aggressive behaviors, RIFT generates traffic flows that preserve the realism of human driving styles, engage the autonomous vehicle in genuine interactive behaviors, and maintain feasibility by ensuring that the resulting scenes, though diverse and stress-inducing, remain solvable for the end-to-end method. This balance enables RIFT to evaluate robustness under credible conditions while avoiding degenerate or unsolvable scenarios.
Figure˜5 presents representative interactions between RIFT-generated traffic flows and two representative end-to-end driving models, UniAD and VAD. As shown, RIFT adapts seamlessly to different driving policies, producing realistic and interactive scenes where the autonomous vehicle must negotiate with surrounding traffic. These results underscore the capability of RIFT to provide realistic yet interactive closed-loop evaluations, highlighting its potential as a versatile tool for testing the robustness of end-to-end AV systems.
Figure 5: Representative closed-loop interactions between RIFT-generated traffic flows and end-to-end autonomous driving algorithms. UniAD (top) and VAD (bottom) are shown interacting with surrounding vehicles orchestrated by RIFT, which preserves realistic driving styles while enabling dynamic CBV–AV interactions. The controlled background vehicle (CBV) is highlighted in purple, the autonomous vehicle (AV, end-to-end) in red, and other background vehicles (BVs) in blue.
Appendix D Additional Results
D.1 Detailed Qualitative Results of Style-Level Controllability
As discussed in Section˜5.4, we investigate the style-level controllability of RIFT under different reward configurations. The aggressive variant applies a reduced collision penalty and places greater emphasis on driving efficiency (Table˜6), encouraging assertive behaviors such as overtaking. In contrast, the normal configuration imposes a higher collision penalty to promote safer and more conservative driving behaviors.
Quantitative results in Table˜3 show that the aggressive variant achieves greater driving efficiency at the expense of more frequent collisions and off-road events. To complement these findings, Figure˜6 presents a qualitative comparison in a single-lane intersection scenario where a leading BV halts at a stop sign. The aggressive CBV variant attempts an overtaking maneuver, resulting in a collision, whereas the normal CBV variant yields and waits, demonstrating distinct behavioral patterns induced by different reward preferences. These results highlight the controllability of RIFT in modulating driving style according to user-specified reward configuration.
Figure 6: Qualitative illustration of RIFT’s style-level controllability under different reward configurations. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
D.2 Detailed Analysis in Driving Comfort
Metrics. To further evaluate the driving comfort of different CBV methods, we define several comfort metrics based on Bench2Drive, which assesses agent comfort through acceleration and jerk profiles. Specifically, we measure comfort using the following metrics:
  • Uncomfortable Rate (UCR): the percentage of simulation time during which CBVs experience discomfort.
  • Driving Jerk (Jerk): the time derivative of acceleration, quantifying the abruptness of acceleration changes and the smoothness of CBV rollouts.
To determine whether a CBV’s current state is considered comfortable, we adopt the Frame Variable Smoothness (FVS) criterion from Bench2Drive:
[…]
The smoothness variables include longitudinal acceleration (expert bounds: [-4.05, 2.40]), maximum absolute lateral acceleration (expert bounds: [-4.89, 4.89]), and maximum jerk magnitude (expert bounds: [-8.37, 8.37]).
Table 7: Comparison of CBV Comfort Metrics across Various AV Methods. Each metric is evaluated across three random seeds.
Method PDM-Lite PlanT
UCR $\downarrow$ Jerk $\downarrow$ UCR $\downarrow$ Jerk $\downarrow$
Pluto 56.45 $\pm$ 4.14 -0.16 $\pm$ 3.72 50.26 $\pm$ 2.17 -0.42 $\pm$ 3.38
PPO 74.76 $\pm$ 2.71 -0.51 $\pm$ 4.61 74.90 $\pm$ 1.21 0.40 $\pm$ 4.83
FREA 72.40 $\pm$ 1.72 0.29 $\pm$ 4.61 73.48 $\pm$ 3.83 -0.15 $\pm$ 4.91
FPPO-RS 68.33 $\pm$ 1.90 -0.07 $\pm$ 3.96 66.67 $\pm$ 0.82 -0.15 $\pm$ 3.95
SFT-Pluto 68.14 $\pm$ 4.91 -0.06 $\pm$ 4.06 59.78 $\pm$ 4.72 -0.11 $\pm$ 4.00
RS-Pluto 70.31 $\pm$ 4.07 0.32 $\pm$ 4.12 65.18 $\pm$ 2.11 -0.16 $\pm$ 4.07
RTR-Pluto 55.58 $\pm$ 4.76 -0.19 $\pm$ 3.37 45.12 $\pm$ 2.66 -0.14 $\pm$ 3.34
PPO-Pluto 58.29 $\pm$ 2.70 -0.32 $\pm$ 3.70 54.85 $\pm$ 5.82 -0.07 $\pm$ 3.40
REINFORCE-Pluto 68.10 $\pm$ 1.22 0.23 $\pm$ 3.96 64.94 $\pm$ 5.36 -0.11 $\pm$ 3.96
GRPO-Pluto 78.58 $\pm$ 0.59 0.22 $\pm$ 4.62 77.13 $\pm$ 0.65 -0.23 $\pm$ 4.58
RIFT-Pluto (ours) 76.90 $\pm$ 2.82 0.59 $\pm$ 4.12 72.41 $\pm$ 4.02 0.21 $\pm$ 4.44
Figure 7: Controller Performance.
Main Results. The quantitative results of the comfort metrics are presented in Table˜7. All CBV methods exhibit notable levels of driving discomfort. Although the more conservative methods identified in Section˜5.2 achieve relatively lower levels of discomfort, a high baseline of discomfort persists between methods.
To investigate the underlying causes of discomfort, we further decouple the planned trajectories from the executed control actions. In CARLA, most CBV methods rely on PID controllers to transform high-level trajectory waypoints into executable driving commands, including throttle, steering, and brake. As shown in Figure˜7, the upper panel illustrates the speed tracking curve, while the middle panel presents the raw throttle signal and corresponding acceleration profile.
Because trajectory generation is performed state-wise, predicting only the immediate next action, the reference states may vary discontinuously over time. These discontinuities are amplified by the PID controller, whose binary throttle/brake responses induce abrupt changes in acceleration, ultimately leading to discomfort during vehicle operation. Such execution-level instabilities are a major contributor to the discomfort observed across CBV methods.
While many CBV methods attempt to mitigate discomfort through fine-tuning strategies that incorporate post-action feedback via reward shaping or expert action alignment, RIFT adopts a different approach. It employs a state-wise reward model (see Section˜B.7) that quantifies comfort within the trajectory’s virtual forward simulation.
To further analyze this, we visualize both the actual acceleration after executing a selected trajectory and the corresponding virtual-time acceleration (shown in Figure˜7). The results reveal that while virtual-time acceleration aligns with actual motion at the beginning of the trajectory, it underestimates acceleration variations in later segments of the trajectory. This leads to an overly conservative estimation of trajectory-level discomfort, resulting in insufficient supervision during training and reflected in RIFT’s comfort performance in Table˜7.
In summary, the discomfort exhibited by CBV methods can be attributed to two primary sources:
  • Tracking instability, caused by discontinuities in planned trajectories and the limited control fidelity of PID controllers. Discrete, state-wise planning combined with low-resolution, often binary control outputs amplifies acceleration fluctuations and leads to uncomfortable motion.
  • Inadequate comfort modeling, particularly in state-wise reward formulations such as that adopted by RIFT and GPRO. These formulations fail to capture long-term trajectory-level discomfort, leading to insufficient supervision during training and suboptimal comfort performance.
D.3 Visualization of the AV-Centric Closed-Loop Simulation
To qualitatively evaluate the robustness of RIFT across diverse AV-centric scenarios, we provide additional temporal visualizations of closed-loop simulations. As shown in Figure˜8, the traffic scene consists of the autonomous vehicle (AV, controlled by PDM-Lite), background vehicles (BVs), and critical background vehicles (CBVs), which interact dynamically over time.
The visualizations demonstrate the ability of RIFT to generate temporally coherent, realistic, and controllable trajectories across a variety of traffic situations. Even under complex and evolving closed-loop conditions, RIFT maintains stable multimodal behavior, highlighting its effectiveness in simulating realistic and controllable traffic scenarios around the AV.
Appendix E Discussion and Broader Implications
E.1 Use of Large Language Models (LLMs)
The large language model (LLM) was employed as a general-purpose writing assistant during the preparation of this manuscript. Its use was limited to:
  • Language refinement: improving grammar, syntax, and overall readability to ensure clarity and professionalism.
  • Style adjustments: suggesting more concise and precise phrasing while preserving the original meaning and technical content.
The LLM was not involved in research ideation, experimental design, data collection, analysis, or interpretation of results. All intellectual contributions and scientific conclusions are solely those of the authors. This disclosure is provided in accordance with the conference guidelines on LLM usage.
E.2 Limitations and Future Work.
In the current framework, the generation head is frozen during RL fine-tuning, and its reliability stems from the robust trajectory generation capability learned during IL pre-training. However, without reliable expert demonstrations, the realism and robustness of generated trajectories cannot be further improved during RL fine-tuning. This limitation highlights a key avenue for future research: developing methods—potentially leveraging RL or other self-improvement paradigms—that can enhance the trajectory generation quality without relying on expert demonstrations.
E.3 Social Impact
Positive Societal Impacts.
This work presents a practical framework that bridges the gap between realism and controllability in traffic simulation. By decoupling pre-training and fine-tuning, our method enables models pre-trained on real-world datasets to adapt effectively to physics-based simulators, preserving trajectory-level realism and route-level controllability while improving long-horizon closed-loop performance. This paradigm establishes a viable pathway for transitioning data-driven approaches to physics-based simulators, enabling more reliable closed-loop testing and training. Consequently, it advances safer and more robust autonomous systems.
Negative Societal Impacts.
While fine-tuning in physics-based simulators improves closed-loop performance, it may also lead to overfitting to the specific characteristics of the simulator. As a result, the learned policy could struggle to generalize beyond the simulated environment, giving rise to a sim-to-real gap. This gap poses challenges for real-world deployment, as models that perform well in simulation may not retain the same level of reliability when applied to actual autonomous driving systems. Such discrepancies can affect the testing and training stages, highlighting the need for further work to ensure real-world transferability.
(a) Robustness of RIFT across diverse AV-centric traffic scenarios.
(b) Temporal stability of RIFT in closed-loop simulation.
Figure 8: Visualizations of RIFT in diverse AV-centric scenarios. (a) Robustness of RIFT across diverse AV-centric traffic scenarios. (b) Temporal stability of RIFT in closed-loop simulation. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.